### v2

# Energy-Efficient Designs for SIM-Based Broadcast MIMO Systems

Abstract
Stacked intelligent metasurface (SIM), which consists of multiple layers of intelligent metasurfaces, is emerging as a promising solution for future wireless communication systems. In this timely context, we focus on broadcast multiple-input multiple-output (MIMO) systems and aim to characterize their energy efficiency (EE) performance. To explore the potential of SIM, we consider both dirty paper coding (DPC) and linear precoding (LP) and formulate the corresponding EE maximization problems. For DPC, we employ the broadcast channel (BC)- multiple-access channel (MAC) duality to obtain an equivalent problem, and optimize users’ covariance matrices using the successive convex approximation (SCA) and Dinkelbach’s methods. Since the phase shift optimization problem of the SIM meta-elements is one of extremely large size, we adopt a conventional projected gradient-based method for its simplicity. A similar approach is followed for the case of LP. Simulation results show that the proposed optimization methods for the considered SIM-based systems can significantly improve the EE, compared to conventional counterparts. Also, we demonstrate that the number of SIM meta-elements and their distribution across the SIM layers have a significant impact on both the achievable sum-rate and EE performance.
I Introduction
The framework for the future development of International Mobile Telecommunications (IMT) for 2030 highlights sustainability as a fundamental goal for future communication systems, with the aim of minimizing environmental impact through efficient resource usage, reduced power consumption, and lower greenhouse gas emissions [1]. This has made energy-efficient wireless communications a key research focus. At the same time, global mobile network data traffic is projected to reach 563 exabytes (EBs) by 2029 [2], which requires the evolution of network technologies to meet the increasing demand. For example, conventional MIMO systems are advancing toward massive MIMO (mMIMO) and ultra-massive MIMO (umMIMO) systems. However, these technologies face challenges in EE due to the substantial power consumption required to support a large number of radio frequency (RF) chains.
A promising solution to address these challenges is the use of reconfigurable intelligent surfaces, which consist of a large number of programmable metamaterial elements capable of dynamically tuning their electromagnetic (EM) properties. In this way, RISs can modify the incoming waves in a programmable and controllable manner [3]. This capability allows RISs to simultaneously improve multiple performance metrics, such as spectrum efficiency, EE, and coverage. Unfortunately, the significant path-loss of RIS-assisted links greatly limits the potential EE gains from RIS deployment.
To address the limitations of RIS, several innovative metamaterial-based antenna technologies such as holographic radio, dynamic metasurface antennas, and SIMs have been proposed. Holographic radio, also known as holographic MIMO (HMIMO), achieves high directive gain, spectral efficiency (SE) and EE by combining densely packing sub-wavelength metamaterial antenna elements with holographic techniques[4, 5]. Similarly, DMAs consist of multiple microstrips, each composed of a multitude of metamaterial radiating elements and connected to a single RF chain [6]. This design offers better EE than hybrid analog-to-digital (A/D) architectures by eliminating the need for additional power to support numerous phase shifters [7]. However, both HMIMOs and DMAs, as single-layer structures, often require a very large number of elements due to hardware constraints that limit the number of tunable amplitudes/phases.
In contrast, SIMs represent the latest advancement in metamaterial-based antenna technologies. Composed of multiple parallel metasurface layers with nearly passive, programmable meta-elements, SIMs can be integrated with conventional transceivers using a small number of active antennas, offering a low-cost and energy-efficient solution. Inspired by deep neural networks (DNN), SIMs perform signal processing tasks, such as transmit precoding and receive combining, directly in the EM domain when properly controlled and programmed [8]. The close spacing between active antennas and SIM layers significantly reduces the path-loss and eliminates the multiplicative path-loss effects seen in RIS-aided systems. By utilizing whole metasurfaces to achieve array gain and suppress co-channel interference in the EM wave domain, SIMs reduce transmission power requirements while significantly enhancing EE and overall system performance, all with minimal hardware complexity.
Recent studies have investigated various applications of SIMs. In [9], SIMs were exploited to implement a 2D discrete Fourier transform (DFT) for direction of arrival (DOA) estimation. A hybrid channel estimator combining wave and digital domain processing was proposed in [10]. However, it requires perfect orthogonality between user pilot sequences, which is hard to achieve. In [11], the authors minimized the Cramer-Rao bound (CRB) for target estimation in an ISAC system with a SIM. Specifically, using an experimental SIM platform, they showed power variations across layers but without any physical explanation. A general path-loss model for SIM-assisted systems was developed in [12], based on which, algorithms aimed at maximizing the received power were derived. This design can generate complex radiation patterns but requires many layers for discrete phase shifts. Research on sum-rate maximization includes studies on SIM-assisted downlink MISO channels [13] and multi-user systems using statistical channel state information (CSI) [14]. In particular, it was shown that using statistical CSI yields slightly worse performance while achieving lower computational and processing overhead. In [15], deep reinforcement learning (DRL) was shown to outperform conventional optimization approaches for maximizing the sum rate in SIM-assisted communication systems. In uplink SIM-based cell-free MIMO systems, rate optimization using distributed signal processing was studied in [16]. In this scenario, each access point (AP) locally detects user information, which is then combined by the central processing unit (CPU) to recover the final data, though this process can demand substantial link resources in practice.
The integration of SIMs with transmitters and receivers into a SIM-based HMIMO system, capable of performing signal precoding and combining in the wave domain, was proposed in [17]. This system introduced a channel fitting approach that enables the SIM-based HMIMO system to achieve significant channel capacity gains compared to mMIMO and RIS-assisted counterparts. An approach for the mutual information maximization in a SIM-based HMIMO system with discrete signaling was presented in [18], using the cutoff rate as an alternative metric. This study demonstrates that incorporating even a small-scale digital precoder into the system can substantially increase its mutual information.
Despite extensive research on SIM-assisted systems, the EE analysis of SIM-assisted MIMO systems remains largely unexplored. Notably, EE optimization inherently includes SE maximization and transmit power minimization as special cases. Motivated by this gap, we aim to maximize the EE in SIM-aided broadcast systems using DPC and LP. To solve the formulated EE maximization (EEmax) problems, we blend multiple optimization techniques to derive efficient problem-driven algorithms, leveraging the special structure of the considered problems and suitable optimization tools. For DPC, we formulate a joint optimization problem of the covariance matrix of the transmitted signal and the phase shifts of the SIM meta-elements. In the more practical case of LP, we consider a similar joint optimization problem for transmit precoding and SIM phase shifts. For both scenarios, we assume a limited total power budget and impose unit modulus constraints on the SIM meta-elements. Furthermore, perfect CSI is assumed at the BS to explore the full EE potential of SIM-based systems. Note that we do not optimize the number of active transmit antennas, which is left for future work. The main contributions of this paper are summarized as follows:
  • For DPC, we exploit the well-known Gaussian MIMO BC-MAC duality to reformulate the EEmax problem in terms of users’ covariance matrices in the MAC and the phase shifts of the SIM meta-elements. Within the adopted alternating optimization (AO) framework, we present an efficient solution for optimizing the covariance matrices, using a tight concave lower bound of the achievable sum-rate via the SCA method. Dinkelbach’s method is then applied to obtain closed-form expressions for the optimal covariance matrices. Our complexity analysis demonstrates that our proposed method has significantly lower complexity compared to an existing solution [19]. For optimizing the SIM phase shifts, we employ a projected gradient-based method to update all SIM layers in parallel. This approach is viable considering the large size of the considered problem. In this context, we derive closed-form expressions for the complex-valued gradients involved.
  • For LP, we leverage a recent result for the sum-rate maximization that allows for reformulating the considered EEmax problem as an equivalent one, with a greatly reduced dimension. Following this important step, we again invoke the SCA method to derive a quadratic lower bound of the achievable sum-rate and approximate the EEmax problem as a concave fractional program. Next, we apply Dinkelbach’s method to solve the resulting problem, obtaining optimal users’ precoders in closed-form expressions. Similar to the DPC-based scheme, the phase shifts of the SIM meta-elements in this setting are optimized in parallel using a conventional projected gradient-based method.
  • We present efficient implementations of the proposed algorithms, analyze their computational complexity in terms of the number of complex multiplications, and mathematically prove their convergence.
  • We show through simulation results that the proposed algorithms substantially increase the EE in SIM-aided broadcast communication systems, with greater improvements observed for DPC. The results highlight the importance of the proposed precoding schemes in mitigating multi-user interference, especially in systems with a large number of users. We also provide several valuable insights into the design and performance of energy-efficient SIM-based broadcast MIMO systems. First, the EE increases logarithmically with the number of meta-elements per SIM layer, with a nearly double gain when increasing from 49 to 100 compared to 100 to 196. Second, SIM-aided systems with a low number of meta-elements may have lower EE than conventional MIMO systems without SIM integration. Third, a few SIM layers suffice to balance SE gains and power consumption for optimal EE. Fourth, in SIM-aided broadcast systems without digital precoding, optimal EE is achieved by activating a subset of transmit antennas, each transmitting independent data streams. Lastly, at least 3 bits per meta-element are required to keep EE degradation from quantization errors within acceptable limits.
The mathematical notation used in this paper is summarized in Table I.
TABLE I: Notation description.
Symbol Description
$\mathbf{H}$ Matrix
$\mathbf{h}$ Row/column vector
$\mathbb{C}^{m\times n}$ Space of $m\times n$ complex matrices
$\mathbf{H}^{T}$ Transpose of $\mathbf{H}$
$\mathbf{H}^{H}$ Hermitian transpose of $\mathbf{H}$
$|\mathbf{H}|$ Determinant of $\mathbf{H}$
$\operatorname{Tr}(\mathbf{H})$ Trace of $\mathbf{H}$
$\log_{2}(\cdot)$ Binary logarithm
$\ln(\cdot)$ Natural logarithm
$(\cdot)^{+}$ Pseudo-inverse operation
$\left\lfloor\cdot\right\rfloor$ Floor operation
$\left(\cdot\right)^{\ast}$ Complex conjugate
$\left\|\mathbf{H}\right\|$ Frobenius norm
$\operatorname{vec}_{d}(\mathbf{H})$ Vector consisting of the diagonal elements of $\mathbf{H}$
$\mathbf{A}\succeq(\succ)\mathbf{B}$ $\mathbf{A}-\mathbf{B}$ is positive semidefinite (definite)
$\mathbf{I}$ Identity matrix
$\Re(\mathbf{x})$ Real part of $\mathbf{x}$
$\Im(\mathbf{x})$ Imaginary part of $\mathbf{x}$
$\operatorname{diag}(\mathbf{x})$ Matrix with the elements of $\mathbf{x}$ on the main diagonal
$\mathcal{CN}(\cdot,\cdot$) Circularly symmetric complex Gaussian distribution
$|x|$ Modulus of the number $x$
[…]
$\nabla_{\mathbf{x}}f(\cdot)$ Complex gradient of $f(\cdot)$ with respect to (w.r.t.) $\mathbf{x}^{\ast}$
II System Model and Problem Formulation
II-A System Model
System level diagrams of conventional and SIM-aided broadcast systems are shown in Fig. 1. In both cases, we consider a multi-user broadcast system in which a BS with $N_{t}$ transmit antennas communicates with $K$ users, each equipped $N_{r}$ receive antennas. The main difference is in the presence of a SIM at the BS, which comprises of $L$ metasurface layers, each containing $N$ meta-elements. The SIM is controlled by external field programmable gate array (FPGA) devices, which adjust the phase shifts of individual meta-elements for wave domain beamforming. [^1] By considering both SIM (i.e., wave-based precoding) and digital precoding, our system model is general enough to include the wave-based only precoding and digital only precoding as special cases. For the sake of simplicity, we assume that this device is integrated into the BS hardware and communicates with the SIM via a dedicated interface.
The phase shifts of the meta-elements in the $l$-th SIM layer are represented by the diagonal matrix $\boldsymbol{\mathbf{\Phi}}^{l}=\text{diag}(\boldsymbol{\phi}^{l})=\text{diag}([\phi_{1}^{l}\>\phi_{2}^{l}\>\cdots\>\phi_{N}^{l}]^{T})$, where $\phi_{n}^{l}=\exp(j\theta_{n}^{l})$ and $\theta_{n}^{l}$ is the phase shift of the $n$-th element in the $l$-th layer. Signal propagation between two consecutive layers, l1l-1 and $l$, is modeled by the matrix $\mathbf{W}^{l}\in\mathbb{C}^{N\times N}$ for $l=2,3,\dots,L$. Specifically, the signal propagation between the $n$-th meta-element of the (l1)(l-1)-th and the $m$-th meta-element of $l$-th layer of the SIM is presented by[20, Eq. (1)]
$\mathbf{W}^{l}]_{m,n}=\frac{A_{k}\cos\chi_{m,n}}{d_{m,n}}\Bigl(\frac{1}{2\pi d_{m,n}}-\frac{j}{\lambda}\Bigr)e^{j\frac{2\pi d_{m,n}}{\lambda}}$ (1)
where $A_{k}$ is the area of each meta-element, $d_{m,n}$ is the distance between the meta-elements of these two layers, $\chi_{m,n}$ is the angle between the propagation direction and the normal to the (l1)(l-1)-th layer, and $\lambda$ is the wavelength. The signal propagation between the transmit antenna array and the first layer of the SIM is represented by the matrix $\mathbf{W}^{1}\in\mathbb{C}^{N\times N_{t}}$, whose elements can be calculated as in (1).
(a) Conventional broadcast system.
(b) SIM-aided broadcast system.
Figure 1: System level diagrams of conventional and SIM-aided broadcast systems.
The end-to-end channel matrix between the BS and the $k$-th user is expressed as
$\mathbf{H}_{k}=\mathbf{G}_{k}\mathbf{B}\in\mathbb{C}^{N_{r}\times N_{t}}$ (2)
where $\mathbf{G}_{k}$ $\in\mathbb{C}^{N_{r}\times N}$ denotes the channel matrix between the final layer of the SIM and user $k$, and $\mathbf{B}=\boldsymbol{\mathbf{\Phi}}^{L}\mathbf{W}^{L}\cdots\boldsymbol{\mathbf{\Phi}}^{2}\mathbf{W}^{2}\boldsymbol{\mathbf{\Phi}}^{1}\mathbf{W}^{1}\in\mathbb{C}^{N\times N_{t}}$ is the EM response of the transmit SIM. We assume that $\mathbf{G}_{k}$ is perfectly known to the BS to investigate its full theoretical potential.
II-B Dirty Paper Coding (DPC)
To evaluate an upper bound on achievable EE, we first consider DPC, which achieves the capacity region of a Gaussian MIMO BC. The received signal at user $k$ is given by
$\mathbf{y}_{k}=\mathbf{H}_{k}\mathbf{s}_{k}+\sum\nolimits_{j<k}\mathbf{H}_{k}\mathbf{s}_{j}+\sum\nolimits_{j>k}\mathbf{H}_{k}\mathbf{s}_{j}+\mathbf{n}_{k}$ (3)
where $\mathbf{H}_{k}\in\mathbb{C}^{N_{r}\times N_{t}}$ is the channel matrix for user $k$, $\mathbf{s}_{k}\in\mathbb{C}^{N_{t}\times 1}$ is the transmitted signal for user $k$. The noise vector $\mathbf{n}_{k}\in\mathbb{C}^{N_{r}\times 1}$ consists of independent and identically distributed (i.i.d.) elements that are distributed according to $\mathcal{CN}(0,\sigma^{2})$, where $\sigma^{2}$ is the noise variance.
The BS designs $\mathbf{s}_{k}$ with full (non-causal) knowledge of the interference term $\sum_{j<k}\mathbf{H}_{k}\mathbf{s}_{j}$, which results from transmissions to users $1,2,\dots,k-1$. Therefore, DPC can eliminate this interference from previously encoded users, and thus, only the interference caused by users $k+1,\dots,K$ remains. Based on this, the achievable rate of user $k$ is given by
$R_{\text{BC},k}(\mathbf{Q},\boldsymbol{\phi})=\ln\frac{\Bigl|\mathbf{I}+\mathbf{H}_{k}\bigl(\sum_{j\geq k}\mathbf{Q}_{j}\bigr)\mathbf{H}_{k}^{H}\Bigr|}{\Bigl|\mathbf{I}+\mathbf{H}_{k}\bigl(\sum_{j>k}\mathbf{Q}_{j}\bigr)\mathbf{H}_{k}^{H}\Bigr|},$ (4)
where, by slight abuse of notation, $\mathbf{H}_{k}$ stands for $\mathbf{H}_{k}/\sigma$ (i.e., $\mathbf{H}_{k}$ is normalized by the square root of the noise power), $\mathbf{Q}_{k}=\mathbb{E}\bigl\{\mathbf{s}_{k}\mathbf{s}_{k}^{H}\bigr\}\succeq\mathbf{0}$ is the input covariance matrix of user $k$, $\mathbf{Q}=[\mathbf{Q}_{1},\mathbf{Q}_{2},\dots,\mathbf{Q}_{K}]$ and $\boldsymbol{\phi}=[(\boldsymbol{\phi}^{1})^{T},(\boldsymbol{\phi}^{2})^{T},\dots,(\boldsymbol{\phi}^{L})^{T}]^{T}$. These covariance matrices are constrained by the total power budget as
$\sum\nolimits_{k=1}^{K}\operatorname{Tr}(\mathbf{Q}_{k})\leq P_{\max}$ (5)
where $P_{\max}$ is the available transmit power budget.
II-C Multi-user MIMO with LP
Although DPC is a capacity achieving scheme, it has high complexity due to its nonlinear coding nature. On the other hand, LP is much simpler to implement in practice. For LP, the transmitted signal is expressed as
$\mathbf{x}=\sum\nolimits_{k=1}^{K}\mathbf{P}_{k}\mathbf{s}_{k}$ (6)
where $\mathbf{s}_{k}$ is the signal intended for user $k$ and $\mathbf{P}_{k}\in\mathbb{C}^{N_{t}\times N_{r}}$ is the corresponding linear precoder. Thus, the received signal at user $k$ is given by
$\mathbf{y}_{k}=\mathbf{H}_{k}\mathbf{x}+\mathbf{n}_{k}=\mathbf{H}_{k}\mathbf{P}_{k}\mathbf{s}_{k}+\sum_{j=1,j\neq k}^{K}\mathbf{H}_{k}\mathbf{P}_{j}\mathbf{s}_{j}+\mathbf{n}_{k}.$ (7)
By treating the multiuser interference as Gausian noise, the achievable rate of user $k$ is given by
$\!\!\!R_{\text{L},k}(\mathbf{P},\boldsymbol{\phi})=$ $\ln\bigg|\mathbf{I}+\mathbf{H}_{k}\mathbf{P}_{k}\mathbf{P}_{k}^{H}\mathbf{H}_{k}^{H}$
$\times\left(\mathbf{I}+\sum\nolimits_{j=1,j\neq k}^{K}\mathbf{H}_{k}\mathbf{P}_{j}\mathbf{P}_{j}^{H}\mathbf{H}_{k}^{H}\right)^{-1}\bigg|$ (8)
where $\mathbf{H}_{k}=\mathbf{H}_{k}/\sigma$ and the precoding matrices $\mathbf{P}=[\mathbf{P}_{1},\mathbf{P}_{2},\dots,\mathbf{P}_{K}]$ have to satisfy the total power constraint:
$\sum\nolimits_{k=1}^{K}\operatorname{Tr}(\mathbf{P}_{k}\mathbf{P}_{k}^{H})\leq P_{\max}.$ (9)
Note that the achievable rate of user $k$ can alternatively be expressed using covariance matrices by defining $\mathbf{Q}_{k}^{\prime}=\mathbf{P}_{k}\mathbf{P}_{k}^{H}$. However, $\mathbf{Q}_{k}^{\prime}$ differs from $\mathbf{Q}_{k}$ for DPC because of differences in the rate expressions. Some existing solutions impose $\operatorname{rank}(\mathbf{Q}_{k})=N_{r}$ to derive $\mathbf{P}_{k}$, but we do not adopt this approach because optimizing $\{\mathbf{Q}_{k}^{\prime}\}$ introduces additional computational complexity compared to directly optimizing $\{\mathbf{P}_{k}\}$.
II-D Problem Formulation
The objective is to maximize the EE of the considered communication system, which is defined as the ratio of the sum-rate and the total power consumption. The total power consumption is modeled as $P_{\mathrm{tot}}=P_{t}+N_{t}P_{c}+P_{0}+LNP_{s},$ where
$P_{t}=\begin{cases}\sum\nolimits_{k=1}^{K}\operatorname{Tr}(\mathbf{Q}_{k})&\text{DPC}\\ \sum\nolimits_{k=1}^{K}\operatorname{Tr}(\mathbf{P}_{k}\mathbf{P}_{k}^{H})&\text{LP}\end{cases}$ (10)
[…]
For the DPC-based scheme, the EE maximization (EEmax) problem is stated as
$\underset{\mathbf{Q},\boldsymbol{\phi}}{\max}$ $\ \eta_{\mathrm{dpc}}=\frac{W\sum_{k=1}^{K}R_{\text{BC},k}(\mathbf{Q},\boldsymbol{\phi})}{\sum_{k=1}^{K}\operatorname{Tr}(\mathbf{Q}_{k})+N_{t}P_{c}+P_{0}+LNP_{s}}$ (11a)
$\mathrm{s.t.}$ $\ \left|\phi_{n}^{l}\right|=1,\forall l,n$ (11b)
$\ \sum\nolimits_{k=1}^{K}\operatorname{Tr}(\mathbf{Q}_{k})\leq P_{\text{max}};\mathbf{Q}_{k}\succeq\mathbf{0},\forall k,$ (11c)
where $W$ is the system bandwidth. Similarly, the EEmax problem with LP is written as
$\underset{\mathbf{P},\boldsymbol{\phi}}{\max}$ $\ \eta_{\mathrm{lp}}=\frac{W\sum_{k=1}^{K}R_{\text{L},k}(\mathbf{P},\boldsymbol{\phi})}{\sum_{k=1}^{K}\operatorname{Tr}(\mathbf{P}_{k}\mathbf{P}_{k}^{H})+N_{t}P_{c}+P_{0}+LNP_{s}}$ (12a)
$\mathrm{s.t.}$ $\ \left|\phi_{n}^{l}\right|=1,\forall l,n$ (12b)
$\ \sum\nolimits_{k=1}^{K}\operatorname{Tr}(\mathbf{P}_{k}\mathbf{P}_{k}^{H})\leq P_{\text{max}}.$ (12c)
Note that the equality constraints in (11b) and (12b) are treated element-wise. Since $W$ is constant, it is dropped when solving (11) and (12), but included in simulation results in Section VII. Next, we present our proposed methods for solving the above two EEmax problems.
III Proposed Solution to DPC-based SIM
To solve (11), we present an iterative optimization algorithm that alternates between optimizing the covariance matrices and the SIM phase shifts, a prevailing method in SIM-related studies. In particular, for fixed phase shifts, we propose a novel method to optimize the covariance matrices in parallel, using a closed-form water-filling solution, which allocates more transmit power to “stronger” user channels. This nice result is achieved by applying Dinkelbach’s method to iteratively maximize a quadratic lower bound of the objective function. The phase shifts of the SIM meta-elements are optimized by a gradient-based optimization method, which is a viable choice, considering the extremely large size of the SIM.
III-A Covariance Matrix Optimization
For ease of readability, we briefly outlined the steps for optimizing covariance matrices as follows:
  • Reformulate (11) using the BCs-MACs duality to obtain a more tractable equivalent formulation.
  • Derive a quadratic lower bound for the achievable sum-rate under the SCA framework.
  • Apply Dinkelbach’s method to solve the fractional EE objective, leading to a water-filling solution for the covariance matrices.
III-A1 Problem Reformulation
The EEmax problem in (11a) is challenging to solve since its objective function is neither convex nor concave with respect to the optimization variables. To deal with this, we exploit the well-known duality between BCs and MACs [21], which states that the capacity region of the MIMO BC is equal to the capacity achieving region of the dual MIMO MAC. For single-user MIMO systems, this duality means that the capacity of a downlink channel $\mathbf{H}$ equals that of a dual uplink channel $\mathbf{H}^{H}$ under the same power constraint, which is easily understood. Extending this idea to multi-user MIMO systems, for a given channel between the BS and user $k$, the BC channel $\mathbf{H}_{k}$ corresponds to the dual MAC channel $\mathbf{H}_{k}^{H}$. Similarly, each transmit covariance matrix $\mathbf{Q}_{k}$ in the BC has as its counterpart a corresponding dual covariance matrix $\mathbf{S}_{k}$ in the dual MAC. [^3] It should be noted that $\mathbf{Q}_{k}$ can not be directly mapped to $\mathbf{S}_{k}$. Instead, it is required to compute first $\mathbf{S}_{K},\mathbf{S}_{K-1},\dots,\mathbf{S}_{k+1}$ before obtaining $\mathbf{S}_{k}$ [21]. Accordingly, (11) is equivalent to the EEmax problem in the dual MAC, which is expressed as
$\underset{\mathbf{S}}{\operatorname{maximize}}$ $\quad\frac{\ln\left|\mathbf{I}+\sum_{k=1}^{K}\mathbf{H}_{k}^{H}\mathbf{S}_{k}\mathbf{H}_{k}\right|}{\sum_{k=1}^{K}\operatorname{Tr}(\mathbf{S}_{k})+N_{t}P_{c}+P_{0}+LNP_{s}}$ (13a)
$\quad\sum\nolimits_{k=1}^{K}\operatorname{Tr}(\mathbf{S}_{k})\leq P_{\text{max}};\mathbf{S}_{k}\succeq\mathbf{0},\forall k,$ (13b)
where we have omitted (11b) as the phase shifts are fixed, and the power constraint (13b) has the same maximum power as (11c). To solve (13) efficiently, we temporarily drop the power constraint (11c), which results in
$\!\!\!\text{$\underset{\mathbf{S}_{k}\succeq 0}{\operatorname{maximize}}$ }g(\mathbf{S})=$ $\frac{\ln\left|\mathbf{I}+\sum_{k=1}^{K}\mathbf{H}_{k}^{H}\mathbf{S}_{k}\mathbf{H}_{k}\right|}{\sum_{k=1}^{K}\operatorname{Tr}(\mathbf{S}_{k})+N_{t}P_{c}+P_{0}+LNP_{s}}.$ (14)
To appreciate the novelty of our proposed method, we briefly describe the block-coordinate method proposed in [19], which optimizes each $\mathbf{S}_{k}$ sequentially, while other variables are fixed. More precisely, let $\mathbf{S}^{(n)}=[\mathbf{S}_{1}^{(n)},\mathbf{S}_{2}^{(n)},\dots,\mathbf{S}_{K}^{(n)}]$ denote the current iterate. Then the next iterate $\mathbf{S}_{k}^{(n+1)}$ is obtained as
$\mathbf{S}_{k}^{(n+1)}$ $=\arg\underset{\mathbf{S}_{k}\succeq\mathbf{0}}{\max}\>\>g\bigl(\mathbf{S}_{1}^{(n+1)},\ldots,\mathbf{S}_{k-1}^{(n+1)},\mathbf{S}_{k},\mathbf{S}_{k+1}^{(n)},\ldots,\mathbf{S}_{K}^{(n)}\bigr)$
$=\arg\underset{\mathbf{S}_{k}\succeq\mathbf{0}}{\max}\>\>\frac{\ln\left|\mathbf{Z}_{k}\right|+\ln\left|\mathbf{I}+\mathbf{T}_{k}^{H}\mathbf{S}_{k}\mathbf{T}_{k}\right|}{p_{k}+\operatorname{Tr}(\mathbf{S}_{k})}$ (15)
where $\mathbf{Z}_{k}=\mathbf{I}+\sum_{j=1,j\neq k}^{K}\mathbf{H}_{j}^{H}\mathbf{S}_{j}\mathbf{H}_{j}$, $\mathbf{T}_{k}=\mathbf{H}_{k}\mathbf{Z}_{k}^{-1/2}$, and $p_{k}=\sum_{j=1,j\neq k}^{K}\operatorname{Tr}(\mathbf{\mathbf{S}}_{j})+N_{t}P_{c}+P_{0}+LNP_{s}$. Next, to solve (15), the authors of [19] applied Dinkelbach’s method, which leads to the following problem
$\underset{\mathbf{S}_{k}\succeq\mathbf{0}}{\max}F_{\lambda}\bigl(\mathbf{S}\bigr)\triangleq\ln\left|\mathbf{Z}_{k}\right|+\ln\left|\mathbf{I}+\mathbf{T}_{k}^{H}\mathbf{S}_{k}\mathbf{T}_{k}\right|-\lambda_{D}(p_{k}+\operatorname{Tr}(\mathbf{S}_{k}))$ (16)
where $\lambda_{D}$ is a non-negative parameter. For a given $\lambda_{D}$, the above problem can be solved in closed-form [19]. However, such a method requires finding the inverse of $\mathbf{Z}_{k}\in\mathcal{C}^{N_{t}\times N_{t}}$ which has a complexity of $\mathcal{O}(N_{t}^{3})$ in general, and computing the singular value decomposition (SVD) of $\mathbf{T}_{k}$ which has a complexity of $\mathcal{O}(N_{t}^{2}N_{r})$. Thus, the overall complexity of the method presented in [19] is very high.
III-A2 SCA Method: A Tight Lower Bound of Sum Rate
We now propose a more efficient method by leveraging the fact that a stationary solution to (14) is also globally optimal since (14) is a concave-convex program. This observation motivates us to adopt the SCA framework, which is typically used to find a stationary solution for nonconvex programs. To proceed, since $\mathbf{S}_{k}\succeq\mathbf{0}$, we can write $\mathbf{S}_{k}=\mathbf{U}_{k}^{H}\mathbf{U}_{k}$ where $\mathbf{U}_{k}\in\mathbb{C}^{N_{r}\times N_{r}}$. Thus, (14) is equivalent to the following unconstrained optimization problem:
$\text{$\underset{\mathbf{U}_{k}}{\operatorname{maximize}}$ }g(\mathbf{U})=\frac{h(\mathbf{U)}}{\sum_{k=1}^{K}\operatorname{Tr}(\mathbf{U}_{k}^{H}\mathbf{U}_{k})+N_{t}P_{c}+P_{0}+LNP_{s}}.$ (17)
where $h(\mathbf{U)}=\ln\left|\mathbf{I}+\sum_{k=1}^{K}\mathbf{H}_{k}^{H}\mathbf{U}_{k}^{H}\mathbf{U}_{k}\mathbf{H}_{k}\right|$ and $\mathbf{U}=[\mathbf{U}_{1},\mathbf{U}_{2},\dots,\mathbf{U}_{K}]$. It is easy to see that $h(\mathbf{U)}$ can be equivalently rewritten as
$h(\mathbf{U)}=\sum_{j=1}^{K}\ln\Bigl|\mathbf{I}+\Bigl(\mathbf{I}+\sum_{k=j+1}^{K}\mathbf{H}_{k}^{H}\mathbf{U}_{k}^{H}\mathbf{U}_{k}\mathbf{H}_{k}\Bigr)^{-1}\mathbf{H}_{k}^{H}\mathbf{U}_{k}^{H}\mathbf{U}_{k}\mathbf{H}_{k}\Bigr|$
$=\sum_{j=1}^{K}\ln\Bigl|\mathbf{I}+\mathbf{U}_{j}\mathbf{H}_{j}\Bigl(\mathbf{I}+\sum_{k=j+1}^{K}\mathbf{H}_{k}^{H}\mathbf{U}_{k}^{H}\mathbf{U}_{k}\mathbf{H}_{k}\Bigr)^{-1}\mathbf{H}_{j}^{H}\mathbf{U}_{j}^{H}\Bigr|.$
In fact, the $j$-th term in the sum above is the capacity of user $j$ in the dual MAC, using successive interference cancellation [21]. As shown shortly, the above reformulation of $h(\mathbf{U)}$ allows for approximating $h(\mathbf{U)}$ by a “proper bound” to obtain a subproblem that admits a closed-form solution. In this regard, we recall the following inequality [22]:
$\ln\left|\mathbf{I}+\mathbf{V}\mathbf{Y}^{-1}\mathbf{V}^{H}\right|\geq\ln\left|\mathbf{I}+\hat{\mathbf{V}}\hat{\mathbf{Y}}^{-1}\hat{\mathbf{V}}^{H}\right|-\operatorname{Tr}(\hat{\mathbf{V}}\hat{\mathbf{Y}}^{-1}\hat{\mathbf{V}}^{H})$
$+2\mathfrak{R}(\operatorname{Tr}(\hat{\mathbf{V}}\hat{\mathbf{Y}}^{-1}\mathbf{V}^{H}))-\operatorname{Tr}(\mathbf{A}^{H}(\mathbf{V}^{H}\mathbf{V}+\mathbf{Y})),$ (19)
where $\mathbf{A}=\hat{\mathbf{Y}}^{-1}-(\hat{\mathbf{V}}^{H}\hat{\mathbf{V}}+\hat{\mathbf{Y}})^{-1}$. We remark that the above inequality holds for arbitrary $\mathbf{V}$, $\hat{\mathbf{V}}$, $\mathbf{Y}\succ\mathbf{0}$ and $\hat{\mathbf{Y}}\succ\mathbf{0}$, provided that their sizes are compatible and the equality occurs when $\mathbf{V}=\hat{\mathbf{V}}$ and $\mathbf{Y}=\hat{\mathbf{Y}}$. In light of the SCA framework, we denote by $\mathbf{U}_{j}^{(n)}$ the value of $\mathbf{U}_{j}$ after $n$ iterations. Let
Input: $\mathbf{H}$, $\mathbf{S}$, $\lambda^{(0)}\geq 0$, $n\leftarrow 0$
1
2repeat
3
4  for $j=1,2,\cdots,K$ do
5
6    Calculate $\hat{\mathbf{V}}_{j}^{(n)}$, $\hat{\mathbf{Y}}_{j}^{(n)}$, $\mathbf{A}_{j}^{(n)}$, $\mathbf{B}_{j}^{(n)}$
7
8    Decompose $\mathbf{H}_{j}\sum_{l=1}^{j}\mathbf{A}_{l}^{(n)}\mathbf{H}_{j}^{H}=\mathbf{P}_{j}\boldsymbol{\Sigma}_{j}\mathbf{P}_{j}^{H}$
9
10     end for
11
12    Solve for $\lambda_{D}$ in (30)
13
14    Compute $\mathbf{U}_{j}^{(n+1)}$ using (27) for $j=1,2,\cdots,K$
15
16    $\lambda^{(n+1)}\leftarrow\lambda_{D}$
17    $n\leftarrow n+1$
18
19     until $\lambda^{(n)}-\lambda^{(n-1)}>\epsilon_{D}$
20    Calculate all $\mathbf{S}_{k}^{*}=\mathbf{U}_{k}^{(n)H}\mathbf{U}_{k}^{(n)}$
21
22    if $\sum\nolimits_{k=1}^{K}\operatorname{Tr}(\mathbf{S}_{k}^{*})\leq P$ then
23
24      $\mathbf{S}_{\text{opt}}=\mathbf{S}^{*}$
25       else
26
27        $\mathbf{S}_{\text{opt}}=\hat{\mathbf{S}}$ obtained from [23]
28
29         end if
Algorithm 1 Optimization of the covariance matrices in the dual MAC.
$\mathbf{V}=\mathbf{U}_{j}\mathbf{H}_{j}\triangleq\mathbf{V}_{j}\in\mathbb{C}^{N_{r}\times N_{t}},$
$\hat{\mathbf{V}}=\mathbf{U}_{j}^{(n)}\mathbf{H}_{j}\triangleq\hat{\mathbf{V}}_{j}^{(n)}\in\mathbb{C}^{N_{r}\times N_{t}},$
$\mathbf{Y}=\mathbf{I}+\sum_{k=j+1}^{K}\mathbf{H}_{k}^{H}\mathbf{U}_{k}^{H}\mathbf{U}_{k}\mathbf{H}_{k}=\mathbf{I}+\sum_{k=j+1}^{K}\mathbf{V}_{j}^{H}\mathbf{V}_{j}\triangleq\mathbf{Y}_{j}$
and
$\hat{\mathbf{Y}}$ $=\mathbf{I}+\sum\nolimits_{k=j+1}^{K}\mathbf{H}_{k}^{H}\mathbf{U}_{k}^{(n)H}\mathbf{U}_{k}^{(n)}\mathbf{H}_{k}$
$=\mathbf{I}+\sum\nolimits_{k=j+1}^{K}\hat{\mathbf{V}}_{j}^{(n)H}\hat{\mathbf{V}}_{j}^{(n)}\triangleq\hat{\mathbf{Y}}_{j}^{(n)}\in\mathbb{C}^{N_{t}\times N_{t}}.$
Then (19) implies
$h(\mathbf{U})$ $\geq\bar{h}(\mathbf{U};\mathbf{U}^{(n)})=c^{(n)}+\sum\nolimits_{j=1}^{K}2\mathfrak{R}\Bigl(\operatorname{Tr}(\mathbf{B}_{j}^{(n)}\mathbf{U}_{j}^{H})\Bigr)$
$\quad-\sum\nolimits_{j=1}^{K}\operatorname{Tr}(\mathbf{A}_{j}^{(n)H}\sum\nolimits_{k=j}^{K}\mathbf{H}_{k}^{H}\mathbf{U}_{k}^{H}\mathbf{U}_{k}\mathbf{H}_{k}).$ (20)
where
$c^{(n)}=\sum\nolimits_{j=1}^{K}\Bigg[\ln\Bigl|\mathbf{I}+\hat{\mathbf{V}}_{j}^{(n)}\Bigl(\hat{\mathbf{Y}}_{j}^{(n)}\Bigr)^{-1}\hat{\mathbf{V}}_{j}^{(n)H}\Bigr|$
$-\operatorname{Tr}(\hat{\mathbf{V}}_{j}^{(n)}\Bigl(\hat{\mathbf{Y}}_{j}^{(n)}\Bigr)^{-1}\hat{\mathbf{V}}_{j}^{(n)H})-\operatorname{Tr}(\mathbf{A}_{j}^{(n)H})\Bigg]$
$=h(\mathbf{U}^{(n)})-\sum\nolimits_{j=1}^{K}\operatorname{Tr}(\hat{\mathbf{V}}_{j}^{(n)}\Bigl(\hat{\mathbf{Y}}^{(n)}\Bigr)^{-1}\hat{\mathbf{V}}_{j}^{(n)H})-\operatorname{Tr}(\mathbf{A}_{j}^{(n)H})$
$\mathbf{B}_{j}^{(n)}=\bar{\mathbf{B}}_{j}^{(n)}\mathbf{H}_{j}^{H}\in\mathbb{C}^{N_{r}\times N_{r}};\bar{\mathbf{B}}_{j}^{(n)}=\hat{\mathbf{V}}_{j}^{(n)}\Bigl(\hat{\mathbf{Y}}_{j}^{(n)}\Bigr)^{-1}\in\mathbb{C}^{N_{r}\times N_{t}}$
$\mathbf{A}_{j}^{(n)}=\Bigl(\hat{\mathbf{Y}}_{j}^{(n)}\Bigr)^{-1}-\Bigl(\hat{\mathbf{V}}_{j}^{(n)H}\hat{\mathbf{V}}_{j}^{(n)}+\hat{\mathbf{Y}}_{j}^{(n)}\Bigr)^{-1}$
$=\Bigl(\hat{\mathbf{Y}}_{j}^{(n)}\Bigr)^{-1}-\Bigl(\hat{\mathbf{Y}}_{j-1}^{(n)}\Bigr)^{-1}\in\mathbb{C}^{N_{t}\times N_{t}}.$
Regarding the complexity of the above approximation, the following remark is in order.
Remark 1.
Since $\hat{\mathbf{Y}}\in\mathbb{C}^{N_{t}\times N_{t}}$, it may appear that the complexity of constructing the above bound is $\mathcal{O}(N_{t}^{3})$ due to the computation of $\Bigl(\hat{\mathbf{Y}}_{j}^{(n)}\Bigr)^{-1}$. However, we emphasize that this is not the case. Specifically, by invoking the matrix-inversion lemma we can write
$\!\!\!\Bigl(\hat{\mathbf{Y}}_{j}^{(n)}\Bigr)^{-1}$ $=\bigl(\hat{\mathbf{V}}_{j+1}^{(n)H}\hat{\mathbf{V}}_{j+1}^{(n)}+\hat{\mathbf{Y}}_{j+1}^{(n)}\bigr)^{-1}$ (21)
$=\bigl(\hat{\mathbf{Y}}_{j+1}^{(n)}\bigr)^{-1}-\bigl(\hat{\mathbf{Y}}_{j+1}^{(n)}\bigr)^{-1}\hat{\mathbf{V}}_{j+1}^{(n)H}\bigl(\mathbf{I}+$ (22)
$\quad\hat{\mathbf{V}}_{j+1}^{(n)}\bigl(\hat{\mathbf{Y}}_{j+1}^{(n)}\bigr)^{-1}\hat{\mathbf{V}}_{j+1}^{(n)H}\bigr)^{-1}\hat{\mathbf{V}}_{j+1}^{(n)}\bigl(\hat{\mathbf{Y}}_{j+1}^{(n)}\bigr)^{-1}$ (23)
$=\bigl(\hat{\mathbf{Y}}_{j+1}^{(n)}\bigr)^{-1}-\bar{\mathbf{B}}_{j+1}^{(n)H}\bigl(\mathbf{I}+\bar{\mathbf{B}}_{j+1}^{(n)}\hat{\mathbf{V}}_{j+1}^{(n)H}\bigr)^{-1}\bar{\mathbf{B}}_{j+1}^{(n)},$ (24)
for $j=1,2,\ldots,K-1$. The above equation indeed suggests a recursive method to compute $\Bigl(\hat{\mathbf{Y}}_{j}^{(n)}\Bigr)^{-1}$ efficiently. Suppose the inverse of $\hat{\mathbf{Y}}_{j+1}^{(n)}$ is known. Then, we only need to compute the inverse of the matrix $\mathbf{I}+\bar{\mathbf{B}}_{j+1}^{(n)}\hat{\mathbf{V}}_{j+1}^{(n)H}\in\mathbb{C}^{N_{r}\times N_{r}}$, which has complexity of $\mathcal{O}(N_{r}^{3})$, to obtain $\Bigl(\hat{\mathbf{Y}}_{j}^{(n)}\Bigr)^{-1}$. Thus, starting from $\hat{\mathbf{Y}}_{K}^{(n)}=\mathbf{I}$, we can iteratively compute $\Bigl(\hat{\mathbf{Y}}_{j}^{(n)}\Bigr)^{-1}$ for $j=K-1,K-2,\ldots,1$. In this way, $\mathbf{A}_{j}^{(n)}$ is also obtained easily.
Now, using the above lower bound of $h(\mathbf{U})$, we arrive at the following approximate problem
$\text{$\underset{\mathbf{U}_{k}}{\operatorname{maximize}}$ }\frac{\bar{h}(\mathbf{U};\mathbf{U}^{(n)})}{\sum_{k=1}^{K}\operatorname{Tr}(\mathbf{U}_{k}^{H}\mathbf{U}_{k})+N_{t}P_{c}+P_{0}+LNP_{s}}.$ (25)
III-A3 Dinkelbach’s Method: Water-filling Solution
It is important to note that (25) is a concave-convex fractional program since $\bar{h}(\mathbf{U};\mathbf{U}^{(n)})$ is concave. Thus, Dinkelbach’s method can be applied to find the optimal solution, which leads to the following parameterized problem
$\underset{\mathbf{U}_{k}}{\operatorname{maximize}}\;f_{\lambda_{D}}(\mathbf{U})=c^{(n)}+\sum\nolimits_{j=1}^{K}2\mathfrak{R}\Bigl(\operatorname{Tr}(\mathbf{B}_{j}^{(n)}\mathbf{U}_{j}^{H})\Bigr)$
$-\sum\nolimits_{j=1}^{K}\operatorname{Tr}\bigl(\mathbf{A}_{j}^{(n)H}\sum\nolimits_{k=j}^{K}\mathbf{H}_{k}^{H}\mathbf{U}_{k}^{H}\mathbf{U}_{k}\mathbf{H}_{k}\bigr)$
$-\lambda_{D}\Bigl(\sum\nolimits_{j=1}^{K}\operatorname{Tr}(\mathbf{U}_{j}^{H}\mathbf{U}_{j})+N_{t}P_{c}+P_{0}+LNP_{s}\Bigr)$ (26)
where $\lambda_{D}>0$ is a given parameter. Now it is straightforward to see that the above optimization problem can be solved independently for each $\mathbf{U}_{j}$, which admits the following closed-form solution
$\mathbf{U}_{j}=\mathbf{B}_{j}^{(n)}\Bigl(\mathbf{H}_{j}\sum_{l=1}^{j}\mathbf{A}_{l}^{(n)H}\mathbf{H}_{j}^{H}+\lambda_{D}\mathbf{I}\Bigr)^{-1},\>j=1,\ldots K.$ (27)
Substituting (27) into (26) yields
$f_{\lambda_{D}}(\mathbf{U})$ $=\sum_{j=1}^{K}\operatorname{Tr}\Bigl(\mathbf{B}_{j}^{(n)}\Bigl(\mathbf{H}_{j}\sum_{l=1}^{j}\mathbf{A}_{l}^{(n)H}\mathbf{H}_{j}^{H}+\lambda_{D}\mathbf{I}\Bigr)^{-1}\mathbf{B}_{j}^{(n)H}\Bigr)$
$\quad+c^{(n)}-\lambda_{D}\Bigl(N_{t}P_{c}+P_{0}+LNP_{s}\Bigr)$ (28)
Let $\mathbf{H}_{j}\sum_{l=1}^{j}\mathbf{A}_{l}^{(n)H}\mathbf{H}_{j}^{H}=\mathbf{P}_{j}\boldsymbol{\Sigma}_{j}\mathbf{P}_{j}^{H}$ be the eigenvalue decomposition of $\mathbf{H}_{j}\sum_{l=1}^{j}\mathbf{A}_{l}^{(n)H}\mathbf{H}_{j}^{H}$ where $\mathbf{P}_{j}\in\mathbb{C}^{N_{r}\times N_{r}}$ is unitary and $\boldsymbol{\Sigma}_{j}\in\mathbb{C}^{N_{r}\times N_{r}}$ is diagonal. Then, we can rewrite $f_{\lambda_{D}}(\mathbf{U})$ as
$f_{\lambda_{D}}(\mathbf{U})=\sum_{j=1}^{K}\operatorname{Tr}\Bigl(\mathbf{B}_{j}^{(n)}\mathbf{P}_{j}^{H}\Bigl(\boldsymbol{\Sigma}_{j}+\lambda_{D}\mathbf{I}\Bigr)^{-1}\mathbf{P}_{j}\mathbf{B}_{j}^{(n)H}\Bigr)$
$\quad+c^{(n)}+\lambda_{D}\Bigl(N_{t}P_{c}+P_{0}+LNP_{s}\Bigr)$ (29)
$=\sum_{j=1}^{K}\sum_{i=1}^{N_{r}}\frac{d_{ji}}{\sigma_{ji}+\lambda_{D}}+c^{(n)}-\lambda_{D}\Bigl(N_{t}P_{c}+P_{0}+LNP_{s}\Bigr)$ (30)
where $d_{ji}$ is the $i$th diagonal element of $\mathbf{P}_{j}^{H}\mathbf{B}_{j}^{(n)H}\mathbf{B}_{j}^{(n)}\mathbf{P}_{j}$ and $\sigma_{ji}$ is the $i$th diagonal element of $\boldsymbol{\Sigma}_{j}$. Next, we need to find $\lambda_{D}$ such that $f_{\lambda_{D}}(\mathbf{U})=0$, which can be solved using either the bisection method or Newton’s method.
Let $\mathbf{U}_{k}^{\ast}$ be the optimal solution to (25), which is returned by Dinkelbach’s method described above. Then $\mathbf{S}_{k}^{\ast}=(\mathbf{U}_{k}^{\ast})^{{\dagger}}\mathbf{U}_{k}^{\ast}$ is the optimal solution to (14), for $k=1,\ldots,K$. Obviously, if $\sum\nolimits_{k=1}^{K}\operatorname{Tr}(\mathbf{S}_{k}^{\ast})\leq P_{\max}$, then $\mathbf{S}^{\ast}=(\mathbf{S}_{1}^{\ast},\mathbf{S}_{2}^{\ast},\dots,\mathbf{S}_{K}^{\ast})$ is also optimal to (11a). On the other hand, if $\sum\nolimits_{k=1}^{K}\operatorname{Tr}(\mathbf{S}_{k}^{\ast})>P_{\max}$, it is straightforward to see that the optimal solution to (11a) is obtained by solving the sum-rate maximization problem with the sum power constraint, which is defined as:
$\underset{\mathbf{S}_{k}}{\operatorname{maximize}}$ $\ln\left|\mathbf{I}+\sum\nolimits_{k=1}^{K}\mathbf{H}_{k}^{H}\mathbf{S}_{k}\mathbf{H}_{k}\right|$ (31a)
$\operatorname{subject~to}$ $\operatorname{Tr}(\mathbf{S}_{k})\leq P_{\max}.$ (31b)
An efficient method for solving the above problem was proposed in [23], which we omit the details for the sake of brevity. Let $\hat{\mathbf{S}}=(\hat{\mathbf{S}}_{1},\hat{\mathbf{S}}_{2,}\dots,\hat{\mathbf{S}}_{K})$ be the optimal solution to (31). Then it is easy to see that the optimal covariance matrices for (13) are given by
$\mathbf{S}_{\text{opt}}=\begin{cases}\mathbf{S}^{\ast}&\sum\nolimits_{k=1}^{K}\operatorname{Tr}(\mathbf{S}_{k}^{\ast})\leq P_{\max}\\ \hat{\mathbf{S}}&\text{otherwise}.\end{cases}$ (32)
A summary of the described covariance matrix optimization method is presented in Algorithm 1.
III-B SIM Phase Shift Optimization
Since the total power consumption does not depend on the SIM phase shifts, the SIM phase shift optimization problem for fixed $\mathbf{S}$ is formulated as
$\underset{\boldsymbol{\phi}}{\operatorname{maximize}}\quad$ $\kappa(\boldsymbol{\phi})=\ln\left|\mathbf{I}+\sum\nolimits_{k=1}^{K}\mathbf{H}_{k}^{H}\mathbf{S}_{k}\mathbf{H}_{k}\right|$ (33a)
$\operatorname{subject~to}\quad$ $\left|\phi_{n}^{l}\right|=1,\forall l,n.$ (33b)
Considering the large size of the SIM, we adopt a gradient-based method to optimize the phase shifts, which consists of the following iterations:
$\boldsymbol{\phi}^{(n+1)}=\mathcal{P}_{\phi}(\boldsymbol{\phi}^{(n)}+u_{n}\nabla_{\boldsymbol{\phi}}\kappa(\boldsymbol{\phi}^{(n)})),$ (34)
where $\mathcal{P}_{\phi}(\cdot)$ denotes the projection operator that ensures the unit modulus constraint on $\boldsymbol{\phi}$ and $u_{n}$ is the appropriate step size. The gradient of $\kappa(\boldsymbol{\phi})$ w.r.t. the phase shifts of the SIM is determined by the gradients of $\kappa(\boldsymbol{\phi})$ w.r.t. the phase shifts of the constituent SIM layers as
$\nabla_{\boldsymbol{\phi}}\kappa(\boldsymbol{\phi})$ $=[\nabla_{\boldsymbol{\phi}^{1}}\kappa(\boldsymbol{\phi})^{T}\>\cdots\>\nabla_{\boldsymbol{\phi}^{L}}\kappa(\boldsymbol{\phi})^{T}]^{T}.$
The gradient w.r.t. each $\boldsymbol{\phi}^{L}$ is provided in Theorem 1.
Theorem 1.
The gradient of $h(\boldsymbol{\phi})$ w.r.t. $\boldsymbol{\phi}^{L}$ is given by
$\nabla_{\boldsymbol{\phi}^{l}}\kappa(\boldsymbol{\phi})$ $=\operatorname{vec}_{d}(\mathbf{C})$ (35)
where
$\mathbf{D}$ $=\mathbf{I}+\sum\nolimits_{k=1}^{K}\mathbf{H}_{k}^{H}\mathbf{S}_{k}\mathbf{H}_{k}$ (36a)
$\mathbf{C}$ $=\boldsymbol{\Theta}^{l+1:L}\sum\nolimits_{k=1}^{K}\mathbf{G}_{k}^{H}\mathbf{S}_{k}\mathbf{H}_{k}\mathbf{D}^{-1}\boldsymbol{\Theta}^{1:l-1}(\mathbf{W}^{l})^{H}$ (36b)
where $\boldsymbol{\Theta}^{m:n}=(\mathbf{W}^{m})^{H}(\boldsymbol{\mathbf{\Phi}}^{m})^{H}\cdots(\mathbf{W}^{n})^{H}(\boldsymbol{\mathbf{\Phi}}^{n})^{H}$.
Proof:
By differentiating $\kappa(\boldsymbol{\phi})$ w.r.t. $\mathbf{\boldsymbol{\Phi}}^{l}$, we obtain
$\!\!\!\!\!\!\text{d}\kappa(\boldsymbol{\phi})=\operatorname{Tr}\bigg(\mathbf{D}^{-1}\sum\nolimits_{k=1}^{K}\text{d}(\mathbf{H}_{k}^{H}\mathbf{S}_{k}\mathbf{H}_{k})\bigg)=$
$\operatorname{Tr}\bigg(\sum\nolimits_{k=1}^{K}\mathbf{S}_{k}\mathbf{H}_{k}\mathbf{D}^{-1}\text{d}\mathbf{H}_{k}^{H}+\sum\nolimits_{k=1}^{K}\mathbf{D}^{-1}\mathbf{H}_{k}^{H}\mathbf{S}_{k}\text{d}\mathbf{H}_{k}\bigg).$ (37)
Substituting
$\text{d}\mathbf{H}_{k}=\mathbf{G}_{k}\boldsymbol{\mathbf{\Phi}}^{L}\mathbf{W}^{L}\cdots\text{d}\boldsymbol{\mathbf{\Phi}}^{l}\mathbf{W}^{l}\cdots\boldsymbol{\mathbf{\Phi}}^{1}\mathbf{W}^{1}$ (38)
in the previous equation results in
$\text{d}\kappa(\boldsymbol{\phi})=\operatorname{Tr}(\mathbf{C}\text{d}(\mathbf{\boldsymbol{\Phi}}^{l})^{H}+\mathbf{C}^{H}\text{d}\mathbf{\boldsymbol{\Phi}}^{l})$ (39)
where $\mathbf{C}$ is defined in (36b). Hence, we have
$\nabla_{\boldsymbol{\mathbf{\Phi}}^{l}}\kappa(\boldsymbol{\phi})=\mathbf{C}$ (40)
and using $\boldsymbol{\mathbf{\Phi}}^{l}=\text{diag}(\boldsymbol{\phi}^{l}),$ we obtain (35). ∎
Since all the elements of $\boldsymbol{\phi}$ have unit amplitude, the projection $\mathcal{P}_{\phi}(\boldsymbol{\phi})$ is defined by
$\bar{\phi}_{n}^{l}=\begin{cases}\phi_{n}^{l}/|\phi_{n}^{l}|&\phi_{n}^{l}\neq 0,\\ e^{j\alpha},\alpha\in[0,2\pi]&\phi_{n}^{l}=0.\end{cases}$ (41)
[…]
Finally, the EE optimization algorithm for a SIM-aided broadcast system with DPC precoding is given in Algorithm 2.
Input: $\mathbf{H}$, $\mathbf{S}^{(0)}$, $\boldsymbol{\phi}^{(0)}$, $\delta>0$, $u_{0}>0$, $\rho\in(0,1)$, $n\leftarrow 0$
1
2repeat
3
4  Call Algorithm 1 to obtain $\mathbf{S}^{(n+1)}$
5
6  repeat
7
8    $\boldsymbol{\phi}^{(n+1)}=\mathcal{P}_{\phi}(\boldsymbol{\phi}^{(n)}+u_{n}\nabla_{\boldsymbol{\phi}}\kappa(\boldsymbol{\phi}^{(n)}))$
9
[…]
11
12      $u_{n}\leftarrow\rho u_{n}$
13
14       end if
15
[…]
17      Calculate $\mathbf{B}$ and $\mathbf{H}$ for $\boldsymbol{\phi}^{(n+1)}$
18
19      $u_{n+1}\leftarrow u_{n}$
20
21      $n\leftarrow n+1$
22
23       until convergence of $\eta_{\mathrm{dpc}}$ in (11a)
After $\mathbf{S}_{\text{opt}}$ is obtained, optimal covariance matrices in the BC are found using the the BCs-MACs duality.
Algorithm 2 Proposed algorithm for the EE optimization for a SIM-aided broadcast system with DPC precoding.
IV Proposed Solution to LP
In this section, we propose an optimization method for the EEmax problem with LP. Similar to the previous section, the precoding matrices are found in closed-form using Dinkelbach’s method, which also yields a water-filling solution. The phase shifts for the SIM meta-elements are optimized by a gradient-based method.
IV-A Precoding Matrix Optimization
For fixed $\boldsymbol{\phi}$, the precoding matrix optimization in (12) reduces to the EEmax problem in conventional MIMO systems. The structure of our proposed method for precoding matrix optimization as follows:
  • Reduce the dimension of the precoding matrix optimization problem by considering an equivalent reformulation.
  • Apply (19) to obtain a lower bound of the achievable rate for LP.
  • Adopt Dinkelbach’s method to solve the attained problem, which leads to water-filling solution.
IV-A1 Problem Reformulation: Dimension Reduction
It can be observed that the complexity of the direct optimization of $\mathbf{P}$ is proportional to $N_{t}$, which can cause a significant complexity burden even for moderate values of $N_{t}$. Consequently, such a direct optimization method is not practically appealing.
To overcome this issue, we consider an equivalent formulation of (12), introduced in [24], which reduces the problem’s size and computational complexity. Denoting $\mathbf{H}=[\mathbf{H}_{1}^{T}\>\mathbf{H}_{2}^{T}\>\cdots\>\mathbf{H}_{K}^{T}]^{T}\in\mathbb{C}^{KN_{r}\times N_{t}}$, (12) can be equivalently written as
$\underset{\mathbf{X}}{\operatorname{maximize}}$ $f(\mathbf{X})=\frac{\sum_{k=1}^{K}\bar{R}_{k}}{\sum_{k=1}^{K}\operatorname{Tr}(\bar{\mathbf{H}}\mathbf{X}_{k}\mathbf{X}_{k}^{H})+N_{t}P_{c}+P_{0}+LNP_{s}}$ (42a)
$\operatorname{subject~to}$ $\sum\nolimits_{k=1}^{K}\operatorname{Tr}(\bar{\mathbf{H}}\mathbf{X}_{k}\mathbf{X}_{k}^{H})\leq P_{\max},$ (42b)
where
$\bar{R}_{k}=\ln\Bigg|\mathbf{I}+\bar{\mathbf{H}}_{k}\mathbf{X}_{k}\mathbf{X}_{k}^{H}\bar{\mathbf{H}}_{k}^{H}\left(\mathbf{I}+\sum\nolimits_{j=1,j\neq k}^{K}\bar{\mathbf{H}}_{k}\mathbf{X}_{j}\mathbf{X}_{j}^{H}\bar{\mathbf{H}}_{k}^{H}\right)^{-1}\Bigg|,$ (43)
$\mathbf{X}=[\mathbf{X}_{1}\>\mathbf{X}_{2}\>\cdots\>\mathbf{X}_{K}]\in\mathbb{C}^{KN_{r}\times KN_{r}}$ are new optimization variables, and $\bar{\mathbf{H}}_{k}=\mathbf{H}_{k}\mathbf{H}^{H}\in\mathbb{C}^{N_{r}\times KN_{r}}$ is the $k$-th sub-matrix of $\bar{\mathbf{H}}=\mathbf{H}\mathbf{H}^{H}\in\mathbb{C}^{KN_{r}\times KN_{r}}$. The equivalence between (12) and (42) is a result of [24, Prop. 2], and the optimal solutions of the two problems are related as $\mathbf{P}_{k}=\mathbf{H}^{H}\mathbf{X}_{k}$. When comparing the size of $\mathbf{X}$ and $\mathbf{P}=[\mathbf{P}_{1},\mathbf{P}_{2},\dots,\mathbf{P}_{K}]\in\mathbb{C}^{N_{t}\times KN_{r}}$, we see that the equivalent formulation in (42) significantly reduces the number of optimization variables, particularly for systems with large $N_{t}$ (i.e., $N_{t}\gg N_{r}$). Following a similar approach as in Algorithm 1, we first derive a solution for the unconstrained case by dropping (42b), with the solution for the constrained case following immediately.
IV-A2 SCA Method: A Tight Lower Bound of Sum Rate
Upon close inspection, we observe that the denominator of (42) is a quadratic convex function, while the numerator of (42) is neither a convex nor concave function. As a result, solving (42) is nontrivial. To develop an efficient solution, we again exploit the inequality, given in (19), to obtain a lower bound of the achievable rate in (43). Using the identity $\left|\mathbf{I}+\mathbf{Z}_{k}\mathbf{Z}_{k}^{H}\mathbf{Y}_{k}\right|=\left|\mathbf{I}+\mathbf{Z}_{k}^{H}\mathbf{Y}_{k}\mathbf{Z}_{k}\right|$ to reformulate (43), the lower bound of the achievable rate of user $k$ can be expressed as
$\bar{R}_{k}$ $\geq L_{k}=\ln\left|\mathbf{I}+\hat{\mathbf{Z}}_{k}^{H}\hat{\mathbf{Y}}_{k}^{-1}\hat{\mathbf{Z}}_{k}\right|-\operatorname{Tr}(\hat{\mathbf{Z}}_{k}^{H}\hat{\mathbf{Y}}_{k}^{-1}\hat{\mathbf{Z}}_{k})$
$+2\mathfrak{R}(\operatorname{Tr}(\hat{\mathbf{Z}}_{k}^{H}\hat{\mathbf{Y}}_{k}^{-1}\bar{\mathbf{H}}_{k}\mathbf{X}_{k}))$
$-\sum\nolimits_{j=1}^{K}\operatorname{Tr}(\mathbf{X}_{j}^{H}\bar{\mathbf{H}}_{k}^{H}\mathbf{A}_{k}^{H}\bar{\mathbf{H}}_{k}\mathbf{X}_{j})-\operatorname{Tr}(\mathbf{A}_{k}^{H})$ (44)
where $\mathbf{Z}_{k}=\bar{\mathbf{H}}_{k}\mathbf{X}_{k}\in\mathbb{C}^{N_{r}\times N_{r}}$, $\mathbf{\hat{Z}}_{k}=\bar{\mathbf{H}}_{k}\mathbf{X}_{k}^{(n)}\in\mathbb{C}^{N_{r}\times N_{r}}$, $\mathbf{Y}_{k}=\sum\nolimits_{j=1,j\neq k}^{K}\bar{\mathbf{H}}_{k}\mathbf{X}_{j}\mathbf{X}_{j}^{H}\bar{\mathbf{H}}_{k}^{H}+\mathbf{I}\in\mathbb{C}^{N_{r}\times N_{r}}$, $\hat{\mathbf{Y}}_{k}=\sum\nolimits_{j=1,j\neq k}^{K}\bar{\mathbf{H}}_{k}\mathbf{X}_{j}^{(n)}(\mathbf{X}_{j}^{(n)})^{H}\bar{\mathbf{H}}_{k}^{H}+\mathbf{I}\in\mathbb{C}^{N_{r}\times N_{r}}$, $\mathbf{A}_{k}=\hat{\mathbf{Y}}_{k}^{-1}-(\hat{\mathbf{Z}}_{k}\hat{\mathbf{Z}}_{k}^{H}+\hat{\mathbf{Y}}_{k})^{-1}\in\mathbb{C}^{N_{r}\times N_{r}}$, and $\mathbf{X}_{k}^{(n)}$ represents the value of $\mathbf{X}_{k}$ at the $n$-th iteration. Consequently, the resulting approximate problem of (42) is given by
$\underset{\mathbf{X}}{\operatorname{maximize}}$ $\bar{f}(\mathbf{X})=\frac{\sum_{k=1}^{K}L_{k}}{\sum_{k=1}^{K}\operatorname{Tr}(\bar{\mathbf{H}}\mathbf{X}_{k}\mathbf{X}_{k}^{H})+N_{t}P_{c}+P_{0}+LNP_{s}}$ (45)
which is a concave-convex optimization problem.
IV-A3 Dinkelbach’s method: Closed-form Solution
Next, we apply Dinkelbach’s method to solve (45), leading to the following optimization problem
$\underset{\mathbf{X}}{\operatorname{maximize}}\;M_{\lambda_{L}}(\mathbf{X})$ (46)
where
$M_{\lambda_{L}}(\mathbf{X})$ $=\sum\nolimits_{k=1}^{K}L_{k}-\lambda_{L}\Big(\sum\nolimits_{k=1}^{K}\operatorname{Tr}(\bar{\mathbf{H}}\mathbf{X}_{k}\mathbf{X}_{k}^{H})$
$+N_{t}P_{c}+P_{0}+LNP_{s}\Big)$ (47)
and $\lambda_{L}\geq 0$ is a given parameter.
Applying the Karush-Kuhn-Tucker (KKT) first-order optimality condition to (46), we take the gradient w.r.t. $\mathbf{X}_{j}^{*}$ and set it to 0, giving
Input: $\mathbf{H}$, $\mathbf{X}^{(0)}$, $\boldsymbol{\phi}$, $\lambda_{L}^{(0)}\geq 0$, $m\leftarrow 0$
1
2repeat
3
4  $n\leftarrow 0$
5
6  repeat
7
8    Calculate all $\mathbf{\hat{Z}}_{k}$, $\mathbf{\hat{Y}}_{k}$, $\mathbf{A}_{k}$
9
10    Calculate $\mathbf{X}_{\text{opt}}$ according to (49) or (51)
11
12    $\mathbf{X}^{(n+1)}\leftarrow\mathbf{X}_{\text{opt}}$
13
14    $n\leftarrow n+1$
15
16     until $\lambda_{L}^{(m+1)}-\lambda_{L}^{(m)}>\epsilon_{2,L}$
17    $\lambda_{L}^{(m+1)}\leftarrow f(\mathbf{X}^{(n)})$ according to (42a)
18
19    $m\leftarrow m+1$
20
21    $\mathbf{X}^{(0)}\leftarrow\mathbf{X}^{(n)}$
22
23     until $f(\mathbf{X}^{(n+1)})-f(\mathbf{X}^{(n)})>\epsilon_{1,L}$
24    if $\sum_{k=1}^{K}\operatorname{Tr}(\bar{\mathbf{H}}\mathbf{X}_{k}\mathbf{X}_{k}^{H})\leq P_{\max}$ then
25
26      $\mathbf{X}_{\text{opt}}=\mathbf{X}$
27       else
28
29        $\mathbf{X}_{\text{opt}}$ obtained from [24]
30
31         end if
32
33        for $k=1$ $\mathbf{to}$ $K$ do
34
35          $\mathbf{P}_{k}=\mathbf{H}^{H}\mathbf{X}_{\text{opt},k}$
36
37           end for
Algorithm 3 Optimization of the precoding matrices.
$(\hat{\mathbf{Z}}_{j}^{H}\hat{\mathbf{Y}}_{j}^{-1}\bar{\mathbf{H}}_{j})^{H}-\sum\nolimits_{k=1}^{K}\bar{\mathbf{H}}_{k}^{H}\mathbf{A}_{k}^{H}\bar{\mathbf{H}}_{k}\mathbf{X}_{j}-\lambda_{L}\bar{\mathbf{H}}\mathbf{X}_{j}=\boldsymbol{0}$ (48)
To solve for $\mathbf{X}_{j}$ we differentiate two cases. If $\mathbf{H}$ is a row-rank matrix, e.g. (i.e., $KN_{r}<N_{t}$), then $\bar{\mathbf{H}}$ is invertible, and thus (48) results in
$\mathbf{X}_{j}$ $=\left(\sum\nolimits_{k=1}^{K}\bar{\mathbf{H}}_{k}^{H}\mathbf{A}_{k}^{H}\bar{\mathbf{H}}_{k}+\lambda_{L}\bar{\mathbf{H}}\right)^{-1}\bar{\mathbf{H}}_{j}^{H}\hat{\mathbf{Y}}_{j}^{-1}\hat{\mathbf{Z}}_{j}.$ (49)
If $\mathbf{H}$ is column-rank matrix (i.e., $KN_{r}>N_{t}$), we rewrite (48) as
$\mathbf{H}\Bigl(\mathbf{H}_{j}^{H}\hat{\mathbf{Y}}_{j}^{-1}\hat{\mathbf{Z}}_{j}-\sum\nolimits_{k=1}^{K}\mathbf{H}_{k}^{H}\mathbf{A}_{k}^{H}\bar{\mathbf{H}}_{k}\mathbf{X}_{j}-\lambda_{L}\mathbf{H}^{H}\mathbf{X}_{j}\Bigr)=\boldsymbol{0}$ (50)
and finally obtain
$\mathbf{X}_{j}=\Bigl(\sum\nolimits_{k=1}^{K}\mathbf{H}_{k}^{H}\mathbf{A}_{k}^{H}\bar{\mathbf{H}}_{k}+\lambda_{L}\mathbf{H}^{H}\Bigr)^{+}\mathbf{H}_{j}^{H}\hat{\mathbf{Y}}_{j}^{-1}\hat{\mathbf{Z}}_{j}.$ (51)
If the solution obtained from (49) or (51) satisfies the power constraint (42b), then it is also the optimal solution to (42). Otherwise, the optimal $\mathbf{X}$ is determined by solving the achievable sum rate optimization problem
$\underset{\mathbf{X}}{\operatorname{maximize}}$ $\sum\nolimits_{k=1}^{K}\bar{R}_{k}$ (52a)
$\operatorname{subject~to}$ $\sum\nolimits_{k=1}^{K}\operatorname{Tr}(\bar{\mathbf{H}}\mathbf{X}_{k}\mathbf{X}_{k}^{H})\leq P_{\max},$ (52b)
which can be solved by [24, Algorithm 1]. The described optimization algorithm is summarized in Algorithm 3.
IV-B SIM Phase Shift Optimization
For a fixed $\mathbf{P}$, the SIM phase shift optimization problem is formulated as
$\underset{\boldsymbol{\phi}}{\operatorname{maximize}}\quad$ $\tau(\boldsymbol{\phi})=\sum\nolimits_{k=1}^{K}R_{\text{L},k}(\boldsymbol{\phi})$ (53a)
$\operatorname{subject~to}\quad$ $\left|\phi_{n}^{l}\right|=1,\forall l,n,$ (53b)
where we have used the rate expression in (8) instead of (43). This choice not only simplifies the gradient computation of the objective function w.r.t $\boldsymbol{\phi}$ but also leads to faster convergence rate as explained in [25].
The optimization of the phase shifts of the BS SIM follows the same steps as in the case of DPC discussed in subsection III-B. Specifically, the phase shifts are iteratively updated according to
Input: $\mathbf{H}$, $\mathbf{X}^{(0)}$, $\boldsymbol{\phi}^{(0)}$, $\lambda^{(0)}\geq 0$, $\delta>0$, $t_{n}>0$, $\rho\in(0,1)$, $n\leftarrow 0$
1
2repeat
3
4  Call Algorithm 3 to obtain $\mathbf{P}^{(n+1)}$
5
6  repeat
7
8    $\boldsymbol{\phi}^{(n+1)}=\mathcal{P}_{\phi}(\boldsymbol{\phi}^{(n)}+t_{n}\nabla_{\boldsymbol{\phi}}\tau(\boldsymbol{\phi}^{(n)}))$
9
[…]
11
12      $t_{n}\leftarrow\rho t_{n}$
13
14       end if
15
[…]
17      $n\leftarrow n+1$
18
19       until convergence of $\eta_{\mathrm{lp}}$ in (12a)
Algorithm 4 Proposed algorithm for the EE optimization for a SIM-aided broadcast system with LP.
$\boldsymbol{\phi}^{(n+1)}=\mathcal{P}_{\phi}(\boldsymbol{\phi}^{(n)}+t_{n}\nabla_{\boldsymbol{\phi}}\tau(\boldsymbol{\phi}^{(n)})),$ (54)
where $t_{n}$ is the appropriate step size. Also, the gradient $\nabla_{\boldsymbol{\phi}}\tau(\boldsymbol{\phi})$ is given by $\nabla_{\boldsymbol{\phi}}\tau(\boldsymbol{\phi})=[\nabla_{\boldsymbol{\phi}^{1}}\tau(\boldsymbol{\phi})^{T}\>\cdots\>\nabla_{\boldsymbol{\phi}^{L}}\tau(\boldsymbol{\phi}))^{T}]^{T}$, where $\nabla_{\boldsymbol{\phi}^{l}}\tau(\boldsymbol{\phi})$ is expressed in the following theorem.
Theorem 2.
The gradients of $\tau(\boldsymbol{\phi})$ w.r.t. the l-th layer of the SIM at the BS is given by
$\nabla_{\boldsymbol{\phi}^{l}}\tau(\boldsymbol{\phi})=\operatorname{vec}_{d}\left(\sum\nolimits_{k=1}^{K}\boldsymbol{\Theta}^{l+1:L}\mathbf{G}_{k}^{H}\mathbf{F}_{1,k}^{-1}\mathbf{H}_{k}\hat{\mathbf{P}}_{s}\boldsymbol{\Theta}^{1:l-1}\mathbf{W}_{l}^{H}\right)$
$-\operatorname{vec}_{d}\left({}_{k=1}^{K}\sum\nolimits_{k=1}^{K}\boldsymbol{\Theta}^{l+1:L}\mathbf{G}_{k}^{H}\mathbf{F}_{2,k}^{-1}\mathbf{H}_{k}\hat{\mathbf{P}}_{k}\boldsymbol{\Theta}^{1:l-1}\mathbf{W}_{l}^{H}\right)$ (55)
where $\boldsymbol{\Theta}^{m:n}=(\mathbf{W}^{m})^{H}(\boldsymbol{\mathbf{\Phi}}^{m})^{H}\cdots(\mathbf{W}^{n})^{H}(\boldsymbol{\mathbf{\Phi}}^{n})^{H}$, $\hat{\mathbf{P}}_{s}=\sum_{j=1}^{K}\mathbf{P}_{j}\mathbf{P}_{j}^{H}$, $\hat{\mathbf{P}}_{k}=\sum_{j=1,j\neq k}^{K}\mathbf{P}_{j}\mathbf{P}_{j}^{H}$, $\mathbf{F}_{1,k}=\mathbf{I}+\mathbf{H}_{k}\hat{\mathbf{P}}_{s}\mathbf{H}_{k}^{H}$ and $\mathbf{F}_{2,k}=\mathbf{I}+\mathbf{H}_{k}\hat{\mathbf{P}}_{k}\mathbf{H}_{k}^{H}$.
Proof:
See the Appendix. ∎
The outline of the proposed algorithm is given in Algorithm 4.
V Computational Complexity
In this section, the computational complexity for SIM-aided broadcast systems with DPC and LP is obtained by counting the required number of complex multiplications. In the following complexity analysis, for simplicity, we assume that $N,N_{t}\gg N_{r}$ which is a typical case for a SIM-based broadcast communication system.
V-A DPC Precoding
The covariance matrix optimization is performed by Algorithm 1, with the complexity of obtaining $\mathbf{U}$ from $\mathbf{S}$ being negligible. In addition, the complexity of one iteration of inner loop (i.e., lines 2 to 5 in Algorithm 1) is $\mathcal{O}(N_{t}^{2}N_{r})$. Also, the complexity of line 6 is approximately $\mathcal{O}(KN_{t}N_{r}^{2})$, while the complexity of line 7 can be neglected. Let $I_{U}$ denote the number of inner loops (i.e., lines 1 to 1 in Algorithm 1), then the complexity of Algorithm 1 is given by $\mathcal{O}(I_{U}KN_{t}^{2}N_{r})$.
Optimizing the SIM phase shifts requires $\mathcal{O}(KNN_{t}N_{r})$ multiplications for computing $\sum\nolimits_{k=1}^{K}\mathbf{G}_{k}^{H}\mathbf{S}_{k}\mathbf{H}_{k}$. The matrix inversion $\mathbf{A}^{-1}$ and its multiplication contribute $\mathcal{O}(N_{t}^{3}+NN_{t}^{2})$ multiplications. Since $(\mathbf{W}^{m})^{H}(\boldsymbol{\mathbf{\Phi}}^{m})^{H}$ is precomputed and only the diagonal elements are needed in (35), the additional complexity is $\mathcal{O}(LN^{3})$. Hence, the complexity of the gradient calculation is $\mathcal{O}(NN_{t}^{2}+LN^{3})$. After obtaining $\boldsymbol{\phi}^{(n+1)}$, the complexity of calculating $\mathbf{B}$ and $\mathbf{H}$ is $\mathcal{O}(LN^{3})$. In addition, $\mathcal{O}(KN_{t}N_{r}(N_{t}+N_{r})+N_{t}^{3})\approx\mathcal{O}(N_{t}^{3})$ multiplications are needed to obtain $h(\boldsymbol{\phi}^{(n+1)})$. The complexity of the SIM phase shifts optimization is $\mathcal{O}(NN_{t}^{2}+I_{\phi,D}(LN^{3}+N_{t}^{3}))$, where $I_{\phi,D}$ is the number of line search iterations.
Therefore, the overall computational complexity for single iteration of Algorithm 2 is given by
$\!\!\!\!\!\!C_{\text{SIM-DPC}}$ $=\mathcal{O}(I_{U}KN_{t}^{2}N_{r}+NN_{t}^{2}+I_{\phi,D}(LN^{3}+N_{t}^{3})).$ (56)
V-B Linear Precoding (LP)
The optimization of the precoding matrices follows Algorithm 3, where computing $\bar{\mathbf{H}}$ requires $\mathcal{O}(K^{2}N_{t}N_{r}^{2})$ multiplications. Computing the initial $\mathbf{X}$ from $\mathbf{P}$ requires $\mathcal{O}(K^{2}N_{t}N_{r}^{2}+K^{3}N_{r}^{3})$ multiplications. Furthermore, the complexity of calculating all $\mathbf{\hat{Z}}_{k}$, $\mathbf{\hat{Y}}_{k}$ and $\mathbf{A}_{k}$ is $\mathcal{O}(K^{3}N_{r}^{3}+K^{2}N_{r}^{3})$. To determine $\mathbf{X}$,
  • if $N_{t}\geq KN_{r}$, solving (49) requires $\mathcal{O}(K^{3}N_{r}^{3})$ multiplications, or otherwise according to
  • if $N_{t}<KN_{r}$, solving (51) requires $\mathcal{O}(K^{3}N_{r}^{3}+K^{3}N_{t}N_{r}^{2})\approx\mathcal{O}(K^{3}N_{t}N_{r}^{2})$ multiplications.
Thus, the computational complexity of precoding matrix optimization is written as
$C_{x}=\begin{cases}\mathcal{O}(K^{3}N_{r}^{3})&N_{t}\geq KN_{r}\\ \mathcal{O}(K^{3}N_{t}N_{r}^{2})&N_{t}<KN_{r}.\end{cases}$ (57)
The complexity for calculating $\sum_{k}\bar{R}_{k}$ and $f(\mathbf{X}^{(n)})$ can be neglected. Let $I_{X}$ denote the number of inner loops (i.e., lines 3 to 8 in Algorithm 3) , then the complexity of Algorithm 3 is given by $\mathcal{O}(I_{X}C_{x})$.
To optimize the phase shifts of the SIM, we need $\mathcal{O}(KN_{t}^{2})$ multiplications to obtain $\hat{\mathbf{P}}_{s}$ and all $\hat{\mathbf{P}}_{k}$ matrices, while calculating $\mathbf{F}_{1,k}$ and $\mathbf{F}_{2,k}$ requires $\mathcal{O}(KN_{t}N_{r}(N_{t}+N_{r}))$ multiplications, which also holds for $\mathbf{F}_{1,k}^{-1}\mathbf{H}_{k}\hat{\mathbf{P}}_{s}-\mathbf{F}_{2,k}^{-1}\mathbf{H}_{k}\hat{\mathbf{P}}_{k}$. Multiplying these terms with $\mathbf{G}_{k}^{H}$ adds $\mathcal{O}(KNN_{t}^{2})$ multiplications. As mentioned earlier, since $(\mathbf{W}^{m})^{H}(\boldsymbol{\mathbf{\Phi}}^{m})^{H}$ are precomputed and only the diagonal elements are need in (55), the additional complexity is $\mathcal{O}(LN^{3})$. Hence, the complexity of the gradient calculation is $\mathcal{O}(KNN_{t}^{2}+LN^{3})$. After obtaining $\boldsymbol{\phi}^{(n+1)}$, the complexity of calculating $\mathbf{B}$ and $\mathbf{H}$ is $\mathcal{O}(LN^{3})$. To obtain all terms $\sum\nolimits_{j}\mathbf{H}_{k}\mathbf{P}_{j}\mathbf{P}_{j}^{H}\mathbf{H}_{k}^{H},$ we need $\mathcal{O}(KN_{t}N_{r}(N_{t}+N_{r}))$ multiplications. Any additional complexity for computing $g(\boldsymbol{\phi}^{(n+1)})$ can be neglected. Hence, the complexity of optimizing the SIM phase shifts is $\mathcal{O}(KNN_{t}^{2}+I_{\phi,L}LN^{3})$, where $I_{\phi,L}$ is the number of line search steps.
Therefore, the overall computational complexity for one iteration of Algorithm 4 is given by
$C_{\text{SIM-LP}}=\mathcal{O}(I_{X}C_{x}+KNN_{t}^{2}+I_{\phi,L}LN^{3}).$ (58)
VI Convergence
We now prove the convergence of Algorithm 2. First, for a given $\boldsymbol{\phi}$, Algorithm 2 achieves monotonic convergence, which can be shown as follows. From (20), it holds that
$g(\mathbf{U}^{(n+1)})=\frac{h(\mathbf{U}^{(n+1)})}{\sum_{k=1}^{K}\operatorname{Tr}(\mathbf{U}_{k}^{(n+1)H}\mathbf{U}_{k}^{(n+1)})+N_{t}P_{c}+P_{0}+LNP_{s}}$ (59)
$\stackrel{{\scriptstyle\textrm{(a)}}}{{\geq}}\frac{\bar{h}(\mathbf{U}^{(n+1)};\mathbf{U}^{(n)})}{\sum_{k=1}^{K}\operatorname{Tr}(\mathbf{U}_{k}^{(n+1)H}\mathbf{U}_{k}^{(n+1)})+N_{t}P_{c}+P_{0}+LNP_{s}}$ (60)
$\stackrel{{\scriptstyle\textrm{(b)}}}{{\geq}}\frac{\bar{h}(\mathbf{U}^{(n)};\mathbf{U}^{(n)})}{\sum_{k=1}^{K}\operatorname{Tr}(\mathbf{U}_{k}^{(n)H}\mathbf{U}_{k}^{(n)})+N_{t}P_{c}+P_{0}+LNP_{s}}$ (61)
$\stackrel{{\scriptstyle\textrm{(c)}}}{{\geq}}\frac{h(\mathbf{U}^{(n)})}{\sum_{k=1}^{K}\operatorname{Tr}(\mathbf{U}_{k}^{(n)H}\mathbf{U}_{k}^{(n)})+N_{t}P_{c}+P_{0}+LNP_{s}}=g(\mathbf{U}^{(n)})$ (62)
where (a) follows from (20), (b) holds because $\mathbf{U}^{(n+1)}$ is the optimal solution to (25) and that the optimal objective is no less than the objective of a feasible point, and (c) is true because it is easy to check that $\bar{h}(\mathbf{U}^{(n)};\mathbf{U}^{(n)})=h(\mathbf{U}^{(n)})$, i.e. the equality in (20) occurs when $\mathbf{U}=\mathbf{U}^{(n)}$. Regarding (b) above, note again that since (25) is a concave-convex fractional program, Dinkelbach’s method is guaranteed to converge to its optimal solution. Consequently, the sequence $\{g(\mathbf{U}^{(n)})\}$ increases monotonically to an optimal solution to (17), and thus, Algorithm 2 is able to compute an optimal solution to (13). Next, for given covariance matrices, the SIM phase shift is optimized by a standard projected gradient method, for which the convergence is guaranteed. Also, the projected gradient method always yields an improved solution. In other words, Algorithm 2 generates a non-decreasing objective sequence. Since the feasible sets for the covariance matrices and phase shifts are continuous, the objective sequence produced by Algorithm 2 is guaranteed to converge.
Figure 2: Convergence of the proposed algorithms for different number of SIM layers.
Since Algorithm 4 uses a similar method for the optimization of the precoding matrices, as Algorithm 2 for the optimization of the covariance matrices, we can prove, following the same derivation steps, that for a given $\boldsymbol{\phi}$ Algorithm 4 is guaranteed to provide an optimal solution to (25). In addition, a gradient-based optimization of the SIM phase shifts always increase the objective function. Moreover, the feasible sets for the precoding matrices and phase shifts are continuous, which ensures the convergence of the objective sequence in Algorithm 4.
VII Simulation Results
In this section, we evaluate the EE of the proposed schemes with DPC and LP, which are denoted as SIM-DPC and SIM-LP, respectively, through Monte Carlo simulations and compare them with three benchmark schemes:
  1. 1.
    LP-NoSIM: The first benchmark scheme uses LP without SIM integration, with a total power consumption of $P_{t}+N_{t}P_{c}+P_{0}.$
  2. 2.
    SIM-NoLP: The second scheme, omits digital precoding, feeding data streams directly to transmit antennas while optimizing SIM phase shifts as in Algorithm 4. Let $N_{t,k}$ BS antennas are allocated for transmission to the $k$-th user. Since all $N_{t}$ transmit antennas are active, $N_{t,k}$ equals $\left\lfloor N_{t}/K\right\rfloor$ or $\left\lfloor N_{t}/K\right\rfloor+1$. Thus, each data stream of user $k$ is transmitted by $\left\lfloor N_{t,k}/N_{r}\right\rfloor$ or $\left\lfloor N_{t,k}/N_{r}\right\rfloor+1$ BS antennas.
  3. 3.
    SIM-NoLP-RedRF: The third scheme excludes LP but differs from SIM-NoLP by utilizing a reduced number of transmit antennas (i.e., RF chains). Specifically, it activates $KN_{r}$ transmit antennas (i.e., RF chains), each transmitting an independent data stream, while the remaining transmit antennas remain inactive. As such, the total power consumption of $P_{t}+KN_{r}P_{c}+P_{0}+LNP_{s}$. Note that for the case when $KN_{r}>N_{t}$, SIM-NoLP transmits signal to a subset of randomly chosen users, while LP-NoSIM-RedRF cannot operate. To ensure fair comparison, data for SIM-NoLP and LP-NoSIM-RedRF are scaled by $\sqrt{P_{\max}/N_{t}}$ and $\sqrt{P_{\max}/KN_{r}},$ respectively.
Channel Modeling
The channel matrix between the BS and user $k$ is modeled as $\mathbf{G}_{k}=\bar{\mathbf{G}}_{k}\mathbf{R}_{\text{T}}^{1/2}\in\mathbb{C}^{N_{r}\times L}$, where $\bar{\mathbf{G}}_{k}\in\mathbb{C}^{N_{r}\times L}$ denotes the channel between the last SIM layer and the receiver, distributed as $\mathcal{CN}(0,\beta\mathbf{I})$. The path-loss between the BS and user $k$ is given by $\beta(d_{k})=\beta(d_{0})+10b\log_{10}(d_{k}/d_{0})$, where $\beta(d_{0})=20\log_{10}(4\pi d_{0}/\lambda)$ is the free space path-loss at the reference distance $d_{0}$, $b$ is the path-loss exponent, and $d_{k}$ is the distance between the BS and user $k$. Moreover, $\mathbf{R}_{\text{T}}\in\mathbb{C}^{L\times L}$ is the spatial correlation matrix of the SIM, defined as in [17, Eq. (14), (15)].
System Parameters
In the simulation setup, parameters are set as follows: $\lambda=5\,\text{cm}$ (i.e., $f=6\,\text{GHz}$), $N_{t}=16$, $N_{r}=2$, $W=100\thinspace\text{kHz}$, $b=3.5$, $d_{0}=1\,\text{m}$, $L=4$, $K=4$, $P_{\max}=5\,\mathrm{W}$ and $\sigma^{2}=-110\,\text{dB}$. Unless specified otherwise, the number of meta-elements per SIM layer, $N$, is 100. The BS antennas are placed in a planar array parallel to the $xy$-plane, with the midpoint at $(30\,\mathrm{m},0,0)$ and inter-antenna spacing of $\lambda/2$. SIM layers are also placed parallel to the $xy$-plane, with the midpoint of the $l$-th layer at $(30\,\mathrm{m},0,l\lambda/2)$. Meta-elements, each with dimension $\lambda/2\times\lambda/2$, are uniformly placed in a square grid on each SIM layer. Hence, any mutual coupling between different meta-elements can be neglected. Users’ ULAs are parallel to the $x$-axis, with the midpoint of the $k$-th user’s ULA at $(x_{k},y_{k},z_{k})$. Users’ coordinates are uniformly distributed such that $x_{k}\in[1.6\,\text{m},2\,\text{m}],y_{k}\in[-20\,\text{m},20\,\text{m}],z_{k}\in[80\,\text{m},120\,\text{m}]$. The circuit power per RF chain is $P_{c}=30\text{\thinspace dBm},$ the basic power consumption at the BS is $P_{0}=40\thinspace\text{dBm}$ [7, 26], and the power consumption per SIM meta-element is $P_{s}=10\thinspace\text{dBm}$ [27, 28]. Optimization parameters are set $\epsilon_{D}=10^{-6}$ for DPC, and $\epsilon_{1,L}=\epsilon_{2,L}=10^{-6}$ for LP. Gradient-based methods use an initial step size of 1000, $\rho=1/2$ and $\delta=10^{-3}$. Results are averaged over 200 independent channel realizations.
Initialization
For DPC the initial $\mathbf{S}_{k}$ matrices are randomly generated Hermitian matrices satisfying (13b) with equality. For LP the initial $\mathbf{X}_{k}$ are randomly generated matrices satisfying (42b) with equality. The initial elements of $\mathbf{\boldsymbol{\phi}}$ are randomly chosen from the unit circle.
Figure 3: $E$ versus number meta-elements per SIM layer.
The convergence of the proposed algorithms for different numbers of SIM layers is shown in Fig. 2. In general, we can see that the number of iterations required to converge increases with the number of SIM layers (i.e., meta-elements), consistent with previous observations in [18], where a layer-by-layer approach for SIM optimization was used. Notably, the gap in EE performance between SIM-DPC and SIM-LP remains small for any number of SIM layers, making SIM-LP a practical and efficient solution for SIM-based systems. Moreover, the EE is not monotonic with $L$. For SIM-DPC, the EE at convergence is almost the same for $L=4$ and $L=8$, while for SIM-LP, it is slightly higher when $L=8$. This behavior results from the fact that the total power consumption scales linearly with the number of meta-elements, while the achievable sum-rate tends to be saturated [29, Fig. 4].
Fig. 3 shows the EE for different number meta-elements per SIM layer ($N=25,49,100,196)$. The EE of both proposed and the non-precoding benchmark schemes (i.e., SIM-NoLP and SIM-NoLP-RedRF) increases with $N$ but the increase rate gradually declines due to the logarithmic behavior of the achievable sum-rate at near-full transmit power. Among the non-precoding benchmarks, SIM-NoLP-RedRF achieves higher EE than SIM-NoLP due to reduced power consumption from inactive RF chains. Moreover, SIM-NoLP suffers from increased multi-stream interference as in this case each data stream is transmitted from two different transmit antennas without digital precoding, which potentially worsens as $N$ increases, further enlarging the EE gap between these benchmarks. On the other hand, the saturation of the achievable sum-rate, and consequently the EE, with the further increase of the number of meta-elements results in almost the same enlargement of the EE for the two benchmark schemes. For a small $N$, LP-NoSIM achieves the largest EE, but is eventually inferior as other schemes benefit from an increased number of meta-elements. Finally, SIM-DPC achieves slightly higher EE than SIM-LP, with the gap increasing as $N$ grows.
Figure 4: $E$ (blue solid lines) and achievable sum-rate (red dashed lines) versus number SIM layer for the constant number of 400 meta-elements.
Fig. 4 plots the EE and the achievable sum-rate of the considered system versus the number of SIM layers, while keeping a total of $400$ meta-elements. Importantly, we find that SIM-LP is as good as SIM-DPC for both performance metrics, regardless of the distribution of meta-elements on SIM layers. and neither metric is monotonic with $L$. Specifically, both increase from $L=1$ to $L=2$, which is apparently due to the enhanced beamforming capabilities offered by multi-layer structures. However, with $L$ increases further, both performance metrics significantly decrease, potentially approaching LP-NoSIM. This occurs because fewer meta-elements per layer reduce the beamforming capability of each layer, while inter-layer propagation introduces additional path-loss as the number of SIM layers increases. A similar trend is observed in non-precoding benchmarks (i.e., SIM-NoLP and SIM-NoLP-RedRF), where the EE and the achievable sum-rate also decrease for $L>4$. Among them, SIM-NoLP-RedRF achieves higher EE, despite a lower achievable sum-rate as it operates with fewer RF chains.
Figure 5: The EE versus the number users.
Fig. 5 demonstrates the EE for different number of users, $K$. For a small $K$, SIM-NoLP-RedRF provides the highest EE due to its reduced RF chain (and thus power) usage. However, as $K$ increases, its EE declines since the number of RF chains becomes comparable to other schemes, and the lack of amplitude control of the transmitted signal limits its multi-user interference suppression capability [23]. On the other hand, the EE of SIM-NoLP shows a non-monotonic behavior with $K$. Specifically, its EE slightly improves from $K=2$ to $K=4$ due to the SIM being able to exploit multi-user diversity gains. However, when $K$ further increases, the EE of both SIM-NoLP and SIM-NoLP-RedRF declines. The EE of the proposed SIM-based schemes and LP-NoSIM increase with $K$ since digital precoding can effectively mitigate the multi-user interference, allowing them to exploit the multi-user diversity. Moreover, SIM-DPC achieves slightly higher EE than SIM-LP, and the gap widens as $K$ increases. This observation further supports LP as an efficient, practical solution for SIM-based systems. Overall, we conclude that integrating SIM with digital precoding consistently yield the best EE, except for very small $K$.
Figure 6: $E$ versus the maximum transmit power.
Next, Fig. 6 illustrates how the EE varies with the maximum transmit power. All schemes in consideration exhibit a logarithmic EE increase due to the logarithmic growth of the achievable rate. At low transmit power, the proposed schemes (i.e. SIM-DPC and SIM-LP) achieve the same EE, as do the benchmark schemes. As the transmit power increases, SIM-DPC consistently shows a more noticeable improvement in EE over all other schemes, which is attributed to the superior interference suppression capabilities of DPC. As expected, the proposed schemes outperform benchmark schemes, with, SIM-NoNP-RedRF obtaining the highest EE among them because of fewer RF chains being used. The difference in the EE between SIM-LP and other benchmark schemes increases with transmit power stabilizes at higher transmit power levels.
To assess the impact of realistic imperfections, Fig. 7 presents the EE of the proposed schemes for discrete SIM phase shifts. The EE generally deteriorates as the number of quantization bits decreases, especially for SIMs with a larger number of layers since the quantization errors are amplified. As a rule of thumb, at least 3 bits per meta-element are required for SIMs for a small $L$ (e.g., 2 or 4) to maintain acceptable EE levels, while larger $L$ requires finer resolution.
Figure 7: EE for the case of discrete SIM phase shifts.
Finally, Fig. 8 shows the per-iteration computational complexity of the proposed optimization schemes versus the number of transmit antennas. The LP-NoSIM scheme achieves the lowest complexity compared to other schemes, as precoding without SIM involves much fewer elements than SIM meta-elements. Due to the same reason, the complexity of the benchmark schemes without precoding is very similar to that of the proposed schemes, and it is not even possible to observe any deference between DPC and LP.
Figure 8: Computational complexity versus the number transmit antennas.
Figure 9: Computational complexities of the proposed schemes with DPC and LP that are without the SIM, and the scheme from [19] versus the number transmit antennas.
To better highlight the difference between DPC and LP, we compare the computational complexities of their appropriate schemes without SIM (i.e., DPC-NoSIM and LP-NoSIM) in Fig. 9. As a benchmark, we consider the DPC-based scheme from [19], also without SIM, referred to as DPC-bench. Its complexity is $O(I_{D}K(N_{t}^{2}N_{r}+N_{t}^{3}))$, where $I_{D}$ is the required number of loops. It can be observed that LP-NoSIM has the highest complexity for $N_{t}=6$ and $N_{t}<KN_{r}$. As $N_{t}$ further increases, the complexities of DPC-NoSIM and DPC-bench become larger than that of LP-NoSIM. This is because the LP-NoSIM complexity remains approximately constant for $N_{t}\geq KN_{r}$, as indicated by (57)). In addition, DPC-bench has a higher complexity than DPC-NoSIM, despite both achieving the same energy efficiency.
VIII Conclusion, Insights, and Future Work
In this paper, we studied the EE maximization in a SIM-aided broadcast system, considering both DPC and LP at the BS. For DPC, we exploited the well-known BC-MAC duality to reformulate the optimization problem and employed a SCA-based technique to derive a tight lower bound on the achievable sum-rate, which was then optimized using Dinkelbach’s method. A similar approach was used to optimize precoders in the case of LP. The phase shifts of the SIM meta-elements were optimized using a conventional projected gradient-based method due to its simplicity. Also, we conducted a computation complexity analysis of the proposed optimization algorithms and proved their convergence.
We carried out extensive numerical experiments and the simulation results have lead to the following insights which are useful for designing energy-efficient SIM-based systems:
  • The proposed optimization algorithms significantly improve the EE for SIM-aided MIMO broadcast systems.
  • DPC outperforms LP in terms of EE due to its superior interference cancellation, but the EE improvement remains marginal, making LP a practical alternative with lower complexity.
  • SIM-based systems without digital precoding still suffer from higher multi-user interference, leading to lower EE.
  • Increasing the number of meta-elements improves EE, but with diminishing returns due to increased power consumption and the logarithmic nature of the achievable sum-rate.
  • Optimal meta-element distribution across layers is crucial; simply increasing the number of SIM layers can degrade performance due to inter-layer signal propagation loss.
  • Discrete phase shift quantization negatively impacts EE, especially for systems with more SIM layers.
  • In SIM-aided broadcast systems without digital precoding, optimal energy efficient transmission strategies typically involve a subset of active transmit antennas.
For future work, the above insights immediately motivate investigating the optimal distance between SIM layers to achieve the best EE performance. Another interesting topic is studying the performance of SIMs in a dynamic environment with user mobility, which would involve adapting the proposed optimization algorithms.
[Proof of Theorem 2]
Differentiating $R_{k}(\boldsymbol{\phi})$ in (8) w.r.t. $\mathbf{H}_{k}$ yields
$\text{d}R_{\text{L},k}(\boldsymbol{\phi})=\text{d}\ln\left|\mathbf{F}_{1,k}\right|-\text{d}\ln\left|\mathbf{F}_{2,k}\right|$ (63)
where
$\!\!\!\text{d}\ln\left|\mathbf{F}_{1,k}\right|$ $=\operatorname{Tr}\left(\hat{\mathbf{P}}_{s}\mathbf{H}_{k}^{H}\mathbf{F}_{1,k}^{-1}\text{d}\mathbf{H}_{k}+\mathbf{F}_{1,k}^{-1}\mathbf{H}_{k}\hat{\mathbf{P}}_{s}\text{d}\mathbf{H}_{k}^{H}\right)$ (64)
$\!\!\!\text{d}\ln\left|\mathbf{F}_{2,k}\right|$ $=\operatorname{Tr}\left(\hat{\mathbf{P}}_{s}\mathbf{H}_{k}^{H}\mathbf{F}_{2,k}^{-1}\text{d}\mathbf{H}_{k}+\mathbf{F}_{2,k}^{-1}\mathbf{H}_{k}\hat{\mathbf{P}}_{k}\text{d}\mathbf{H}_{k}^{H}\right).$ (65)
Substituting (38) into the previous expressions, we obtain
$\text{d}\ln\left|\mathbf{F}_{1,k}\right|$ $=\operatorname{Tr}\left(\mathbf{G}_{1}^{H}\text{d}\boldsymbol{\mathbf{\Phi}}^{l}+\mathbf{G}_{1}\text{d}(\boldsymbol{\mathbf{\Phi}}^{l})^{H}\right)$ (66)
$\text{d}\ln\left|\mathbf{F}_{2,k}\right|$ $=\operatorname{Tr}\left(\mathbf{G}_{2}^{H}\text{d}\boldsymbol{\mathbf{\Phi}}^{l}+\mathbf{G}_{2}\text{d}(\boldsymbol{\mathbf{\Phi}}^{l})^{H}\right)$ (67)
where
$\mathbf{G}_{1}=$ $\boldsymbol{\Theta}^{l+1:L}\mathbf{G}_{k}^{H}\mathbf{F}_{1}^{-1}\mathbf{H}_{k}\hat{\mathbf{P}}_{s}\boldsymbol{\Theta}^{1:l-1}\mathbf{W}_{l}^{H}$ (68)
$\mathbf{G}_{2}=$ $\boldsymbol{\Theta}^{l+1:L}\mathbf{G}_{k}^{H}\mathbf{F}_{2}^{-1}\mathbf{H}_{k}\hat{\mathbf{P}}_{k}\boldsymbol{\Theta}^{1:l-1}\mathbf{W}_{l}^{H}.$ (69)
After a few simple mathematical steps, we get
$\nabla_{\boldsymbol{\phi}^{l}}R_{\text{L},k}(\boldsymbol{\phi})=\operatorname{vec}_{d}\left(\boldsymbol{\Theta}^{l+1:L}\mathbf{G}_{k}^{H}\mathbf{F}_{1,k}^{-1}\mathbf{H}_{k}\hat{\mathbf{P}}_{s}\boldsymbol{\Theta}^{1:l-1}\mathbf{W}_{l}^{H}\right)$
$-\operatorname{vec}_{d}\left(\boldsymbol{\Theta}^{l+1:L}\mathbf{G}_{k}^{H}\mathbf{F}_{2,k}^{-1}\mathbf{H}_{k}\hat{\mathbf{P}}_{k}\boldsymbol{\Theta}^{1:l-1}\mathbf{W}_{l}^{H}\right).$ (70)
From this gradient expression for the achievable rate of user $k$, we can easily obtain the appropriate gradients for all other users. After summing all these gradient expressions, we obtain (55). This completes the proof.